### v2

# On the Learning with Augmented Class via Forests

Abstract
Decision trees and forests have achieved successes in various real applications, most working with all testing classes known in training data. In this work, we focus on learning with augmented class via forests, where an augmented class may appear in testing data yet not in training data. We incorporate information of augmented class into trees’ splitting, that is, augmented Gini impurity, a new splitting criterion is introduced to exploit some unlabeled data from testing distribution. We then develop the Learning with Augmented Class via Forests (short for LACForest) approach, which constructs shallow forests according to the augmented Gini impurity and then splits forests with pseudo-labeled augmented instances for better performance. We also develop deep neural forests via an optimization objective based on our augmented Gini impurity, which essentially utilizes the representation power of neural networks for forests. Theoretically, we present the convergence analysis for our augmented Gini impurity, and we finally conduct experiments to evaluate our approaches. The code is available at https://github.com/nju-xuf/LACForest.
1 Introduction
How to handle distribution changes has become an important problem in a non-stationary learning environment Zhou (2022), and recent years have witnessed increasing attentions with various applications Gama et al. (2014); Geng et al. (2021); Wang et al. (2024). This work focuses on learning with augmented class, that is, the class distribution changes and an augmented class unseen in the training data may emerge during the testing process Da et al. (2014). Here, we take object recognition in autonomous driving for example: new objects may appear on roads yet not in the historical labeled data, and a reliable learning system should make good predictions over both known classes and augmented class.
Various approaches have been developed for learning with augmented class. Da et al. Da et al. (2014) studied decision boundaries for augmented class under the low-density assumption. Bendale and Boult Bendale and Boult (2016) and Rudd et al. Rudd et al. (2017) estimated the probability of an instance belonging to augmented class on the basis of extreme value theory. Mendes Júnior et al. Mendes Júnior et al. (2017) detected augmented class by a nearest neighbor method, and Liu et al. Liu et al. (2018) presented PAC guarantees for the detection of augmented class. Some generative networks have also been applied to learn augmented class Ge et al. (2017); Neal et al. (2018); Chen et al. (2021). Zhang et al. Zhang et al. (2020) gave an unbiased risk estimation by exploiting unlabeled data, and Shu et al. Shu et al. (2023) generalized such approach to arbitrary loss functions.
Decision trees and forests have achieved great successes in various applications with strong generalization in handling discrete features and exploring local regions Cutler et al. (2007); Qi (2012); Grinsztajn et al. (2022); Costa and Pedreira (2023). Most previous studies worked on the same distribution between training and testing data. For augmented class, Mu et al. Mu et al. (2017) and Liu et al. Liu et al. (2018) only studied its detection from labeled data of known classes, whereas it remains open on how to construct trees for learning with augmented class by exploiting unlabeled data from testing data.
This work aims to incorporate some useful information of augmented class into the construction of forests, and our main contributions can be summarized as follows:
  • We introduce a new splitting criterion, i.e., augmented Gini impurity, to incorporate information of augmented class from unlabeled data during the trees’ splitting. We develop the LACForest approach for learning with augmented class, which constructs shallow forests by augmented Gini impurity and considers pseudo-labeled augmented instances for better performance.
  • For complex data with intrinsic structures (e.g., images), we develop deep neural forests for learning with augmented class, because of their powerful representations with an end-to-end training manner. We propose a new optimization objective for deep neural forests to learn augmented class, and the basic idea is to extend our augmented Gini impurity into a differentiable form by considering the mechanism of deep neural trees.
  • From a theoretical view, we present the convergence analysis of our augmented Gini impurity with respect to both decision trees and deep neural trees. We finally conduct extensive experiments to validate the effectiveness of our proposed approaches and perform some parameter influence analysis.
The rest of this work is organized as follows: Section 2 gives some preliminaries. Section 3 introduces augmented Gini impurity and LACForest approach. Section 4 presents our deep neural approach. Section 5 conducts some extensive experiments. Section 6 concludes with future works.
2 Preliminaries
Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\{1,\ldots,\kappa,\kappa+1\}$ be the instance and class spaces, respectively. Suppose that $\mathcal{D}$ is an underlying (unknown) distribution over $\mathcal{X}\times\mathcal{Y}$, and denote by $\mathcal{D}_{\mathcal{X}}$ its marginal distribution over $\mathcal{X}$. We focus on learning with augmented class Da et al. (2014), where the first $\kappa$ classes can be observed in training data, while the $(\kappa+1)$-th class will merely emerge in testing data, which is known as an augmented class in training phase.
For known classes and augmented class, let $\mathcal{D}_{\text{kc}}$ and $\mathcal{D}_{\text{ac}}$ be the marginal distributions over space $\mathcal{X}\times\{1,\ldots,\kappa\}$ and $\mathcal{X}\times\{\kappa+1\}$ from distribution $\mathcal{D}$, respectively. We introduce the class shift assumption as follows:
Assumption 1.
We say that distribution $\mathcal{D}$ and its marginal distributions $\mathcal{D}_{\text{kc}}$ and $\mathcal{D}_{\text{ac}}$ satisfy class shift assumption if
$\mathcal{D}=(1-\theta)\mathcal{D}_{\text{kc}}+\theta\mathcal{D}_{\text{ac}}\ \text{ for some constant }\ \theta\in(0,1)\ .$ (1)
This assumption correlates data distributions for known classes and augmented class, which has been well-studied for learning with augmented class Zhang et al. (2020); Shu et al. (2023) and open-set recognition Scheirer et al. (2013).
We focus on the semi-supervised setting for learning with augmented class Da et al. (2014); Liu et al. (2018); Zhang et al. (2020); Shu et al. (2023), and the goal is to learn a function $f\colon\mathcal{X}\rightarrow\mathcal{Y}$ from labeled data $S_{l}$ and unlabeled data $S_{u}$ with
$S_{l}=\{(\bm{x}^{l}_{1},y_{1}),\ldots,(\bm{x}^{l}_{n_{l}},y_{n_{l}})\}\text{ and }S_{u}=\{\bm{x}^{u}_{1},\ldots,\bm{x}^{u}_{n_{u}}\}\ .$
Here, each labeled example $(\bm{x}^{l}_{i},y_{i})$ is drawn i.i.d. from $\mathcal{D}_{\text{kc}}$, and each unlabeled instance $\bm{x}^{u}_{j}$ is sampled i.i.d. from $\mathcal{D}_{\mathcal{X}}$.
Let $\mathbb{I}[\cdot]$ be the indicator function, which returns $1$ if the argument is true and 00 otherwise. For integer $k>0$, we write $[k]=\{1,2,\cdots,k\}$, and denote by $\Delta_{k}$ the $k$-dimensional simplex. For $z\in\mathbb{R}$, we denote by $\lfloor z\rfloor$ the largest integer no more than $z$, and define $(z)_{+}=\max(0,z)$. Let $|A|$ be the cardinality of set $A$.
3 Our LACForest Approach
This section proposes the LACForest approach for learning with augmented class based on random forests, and the core idea is to introduce a new splitting criterion, augmented Gini impurity, to incorporate potential information of augmented instances during the construction of decision trees.
3.1 Augmented Gini Impurity
For instance space $\mathcal{C}\subseteq\mathcal{X}$, unlabeled data $S_{u}$ and labeled data $S_{l}$, we introduce
$S_{\mathcal{C},u}=S_{u}\cap\mathcal{C}\ \text{ and }\ S_{\mathcal{C},l}=\{(\bm{x},y)\colon(\bm{x},y)\in S_{l}\text{ and }\bm{x}\in\mathcal{C}\}.$
Denote by $n_{\mathcal{C},l}=|S_{\mathcal{C},l}|$ and $n_{\mathcal{C},u}=|S_{\mathcal{C},u}|$. We could define the augmented Gini impurity as a splitting criterion.
Definition 1.
For an instance space $\mathcal{C}\subseteq\mathcal{X}$, we define the augmented Gini impurity $\mathcal{G}_{\mathcal{C}}(S_{l},S_{u})$ w.r.t. $S_{l}$ and $S_{u}$ as
$\mathcal{G}_{\mathcal{C}}(S_{l},S_{u})=1-\sum_{k\in[\kappa+1]}\vartheta^{2}_{\mathcal{C},k}(S_{l},S_{u})\ .$ (2)
Here, $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$ is defined as
$\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})=\mathbb{I}[n_{\mathcal{C},u}>0]\Big{(}1-\frac{(1-\theta)n_{u}n_{\mathcal{C},l}}{n_{l}\max(1,n_{\mathcal{C},u})}\Big{)}_{+}\ ,$
where $\theta$ is given in Eqn. (1); and for $k\in[\kappa]$, we define
$\vartheta_{\mathcal{C},k}(S_{l},S_{u})=\sum_{(\bm{x},y)\in S_{l}}(1-\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u}))\frac{\mathbb{I}[\bm{x}\in\mathcal{C},y=k]}{\max(1,n_{\mathcal{C},l})}\ .$
Figure 1: An intuitive illustration on the difference between our criterion and original Gini impurity over a 2-dimensional dataset, by considering augmented class.
This definition essentially follows the original Gini impurity Breiman (1984), except with an additional term $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$, which aims to incorporate some information of augmented class into impurity measure.
Figure 1 presents an intuitive illustration on the difference between our augmented and original Gini impurity over a 2-dimensional synthetic dataset. It is observable that our augmented Gini impurity could properly take augmented data into consideration compared with Gini impurity.
For $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$, we also have
Lemma 1.
For $\delta\in(0,1)$ and instance space $\mathcal{C}\subseteq\mathcal{X}$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
$\left|\Pr_{(\bm{x},y)\in\mathcal{D}}[y=\kappa+1|\bm{x}\in\mathcal{C}]-\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})\right|$
$\leq$ $O\left(\frac{\sqrt{\ln(1/\delta)}}{\gamma}\left(\frac{1}{\sqrt{n_{l}}}+\frac{\sqrt{\gamma}+1}{\sqrt{n_{u}}}\right)\right)\ ,$
if $n_{\mathcal{C},u}/n_{u}\geq\gamma$ for some constant $\gamma\in(0,1)$.
Let $\mathcal{L}^{*}_{\mathcal{C}}$ be the optimal squared loss over distribution $\mathcal{D}$ under the condition $\bm{x}\in\mathcal{C}$ as follows:
$\mathcal{L}^{*}_{\mathcal{C}}=\min_{\bm{w}\in\Delta_{\kappa+1}}E_{(\bm{x},y)\sim\mathcal{D}}\left[\|\bm{w}-\tilde{y}\|_{2}^{2}|\bm{x}\in\mathcal{C}\right]\ ,$ (3)
where $\tilde{\bm{y}}\in\mathbb{R}^{\kappa+1}$ is the one-hot encoding of $y\in[\kappa+1]$. Intuitively, $\mathcal{L}^{*}_{\mathcal{C}}$ shows the minimum expected squared loss under the same prediction for all instances in space $\mathcal{C}$.
Based on Lemma 1 and Assumption 1, we have
Theorem 1.
For $\delta\in(0,1)$ and instance space $\mathcal{C}\subseteq\mathcal{X}$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
$\left|\mathcal{L}^{*}_{\mathcal{C}}-\mathcal{G}_{\mathcal{C}}(S_{l},S_{u})\right|\leq O\left(\frac{\kappa\sqrt{\ln(\kappa/\delta)}}{\gamma/(1+\sqrt{\gamma})}\left(\frac{1}{\sqrt{n_{l}}}+\frac{1}{\sqrt{n_{u}}}\right)\right)$
if $n_{\mathcal{C},u}/n_{u}\geq\gamma$ and $n_{\mathcal{C},l}/n_{l}\geq\gamma$ for constant $\gamma\in(0,1)$.
This theorem shows that our augmented Gini impurity can be viewed as a good estimation of $\mathcal{L}^{*}_{\mathcal{C}}$ with the convergence rate $O(1/\sqrt{\min\{n_{l},n_{u}\}})$. The detailed proof is given in Appendix A.2, which involves the equivalence between splitting criterions and loss functions, as well as the convergence of $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$ in Lemma 1.
3.2 Construction of LACForest
A tree model can be constructed by partitioning instance space into disjoint rectangular nodes recursively. Given node $\mathcal{C}\subseteq\mathcal{X}$, splitting feature $j\in[d]$ and splitting point $a\in\mathbb{R}$, we denote by the left and right children as
$\mathcal{C}^{l}_{j,a}=\{\bm{x}\in\mathcal{C}\colon x_{j}\leq a\}\ \text{ and }\ \mathcal{C}^{r}_{j,a}=\{\bm{x}\in\mathcal{C}\colon x_{j}>a\}\ ,$
respectively. We then solve the optimal splitting feature $j^{*}$ and splitting point $a^{*}$ from the following optimization
$(j^{*},a^{*})\in{\arg\max}_{j,a}\left\{R_{\mathcal{C}}(S_{l},S_{u},j,a)\right\}\ ,$ (4)
where the reduction of augmented Gini impurity is given by
$R_{\mathcal{C}}(S_{l},S_{u},j,a)=\mathcal{G}_{\mathcal{C}}(S_{l},S_{u})\\ -\frac{n_{\mathcal{C}^{l}_{j,a},u}}{n_{\mathcal{C},u}}\mathcal{G}_{\mathcal{C}^{l}_{j,a}}(S_{l},S_{u})-\frac{n_{\mathcal{C}^{r}_{j,a},u}}{n_{\mathcal{C},u}}\mathcal{G}_{\mathcal{C}^{r}_{j,a}}(S_{l},S_{u})\ .$
From Theorem 1, we can distinguish augmented class from known classes in node $\mathcal{C}$ by solving Eqn. (4), as it converges to optimizing the squared loss as training data increases. From a theoretical view, we require sufficient training data in each node $\mathcal{C}$, i.e., $n_{\mathcal{C},l}\geq\gamma n_{l}$ and $n_{\mathcal{C},u}\geq\gamma n_{u}$ for some $\gamma\in(0,1)$ in Theorem 1; therefore, we should construct relatively shallow tree models.
On the other hand, we should take deeper tree models for better performance, and it is well-known that better predictions could be obtained by deepening tree models sufficiently Quinlan (1986); Hastie et al. (2009). Hence, our algorithm includes two main steps as follows:
3.2.1 Step-I: Exploration of Augmented Instances
We preliminarily construct $m$ random trees $\mathcal{T}^{\prime}_{1},\cdots,\mathcal{T}^{\prime}_{m}$ for the exploration of augmented instances from Eqn. (4). For simplicity, we present the detailed construction of tree $\mathcal{T}^{\prime}_{1}$, and consider other trees similarly. We initialize $\mathcal{T}^{\prime}_{1}$ with one root node of instance space $\mathcal{X}$, and recursively repeat a procedure for each leaf node as follows:
  • Select a $\tau$-subset $\mathcal{S}$ from $d$ available features randomly without replacement;
  • Solve the optimal splitting feature $j^{*}\in\mathcal{S}$ and point $a^{*}$ according to Eqn. (4) under the constraints
    $\min(n_{\mathcal{C}^{l}_{j,a},l},n_{\mathcal{C}^{r}_{j,a},l})$$\geq$ $\gamma n_{l}\ ,$ (5)
    $\min(n_{\mathcal{C}^{l}_{j,a},u},n_{\mathcal{C}^{r}_{j,a},u})$$\geq$ $\gamma n_{u}\ .$ (6)
    This ensures sufficient instances in each splitting node from Theorem 1;
  • Split current node into left and right children via the optimal splitting feature $j^{*}$ and position $a^{*}$.
The above procedure is stopped if there is no feasible solution for Eqn. (4) under constraints in Eqns. (5) and (6).
Given random tree $\mathcal{T}^{\prime}_{1}$ with $t^{\prime}$ leaves $\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{t^{\prime}}$, we present the probability of augmented class for $\bm{x}\in\mathcal{X}$ as
$\mathcal{T}^{\prime}_{1,\kappa+1}(\bm{x})=\sum_{j\in[t^{\prime}]}\vartheta_{\mathcal{C}^{\prime}_{j},\kappa+1}(S_{l},S_{u})\mathbb{I}[\bm{x}\in\mathcal{C}^{\prime}_{j}]\ ,$
where $\vartheta_{\mathcal{C}^{\prime}_{j},\kappa+1}(S_{l},S_{u})$ is an unbiased estimator for the proportion of augmented instances in $\mathcal{C}^{\prime}_{j}$, given by Definition 1.
3.2.2 Step-II: Improvement of Prediction Performance
Algorithm 1 Our LACForest approach
Input: Labeled data $S_{l}$, unlabeled data $S_{u}$, proportion $\theta$, hyper-parameters $m,\tau,\gamma$
1: % Step-I: Exploration of Augmented Instances.
2:for $i\in\{1,\cdots,m\}$ do
3:      Grow random tree $\mathcal{T}^{\prime}_{i}$ based on $S_{l}$ and $S_{u}$ according to the reduction of augmented Gini impurity .
4:end for
5: % Step-II: Improvement of Prediction Performance
6:Calculate $\mathcal{LF}^{\prime}_{\kappa+1}(\bm{x})$ for $\bm{x}\in S_{u}$ by Eqn. (7).
7:Label top $n_{p}=\lfloor\theta n_{u}\rfloor$ instances in $S_{u}$ of the highest $\mathcal{LF}^{\prime}_{\kappa+1}(\bm{x})$ and obtain $\tilde{S}$.
8:for $i\in\{1,\cdots,m\}$ do
9:      Split random tree $\mathcal{T}^{\prime}_{i}$ based on $S_{l}\cup\tilde{S}$ according to the reduction of Gini impurity and obtain $\mathcal{T}_{i}$.
10:end for
11:return $\mathcal{LF}(\bm{x})=\underset{k\in[\kappa+1]}{\arg\max}\sum_{i\in[m]}\mathcal{T}_{i,k}(\bm{x})$.
Given $m$ trees $\mathcal{T}^{\prime}_{1},\cdots,\mathcal{T}^{\prime}_{m}$, we calculate the average augmented score of instance $\bm{x}\in\mathcal{X}$ by
$\mathcal{LF}^{\prime}_{\kappa+1}(\bm{x})=\frac{1}{m}\sum_{i\in[m]}\mathcal{T}^{\prime}_{i,\kappa+1}(\bm{x})\ .$ (7)
We then select $\lfloor\theta n_{u}\rfloor$ instances in $S_{u}$ of the highest average augmented scores, and label them with pseudo-labels of augmented class. Without loss of generality, we denote by
$\tilde{S}=\{(\bm{x}^{u}_{1},\kappa+1),(\bm{x}^{u}_{2},\kappa+1),\cdots,(\bm{x}^{u}_{\lfloor\theta n_{u}\rfloor},\kappa+1)\}\ .$
Intuitively, $\tilde{S}$ retains the most likely augmented instances for further partition, and similar approaches have been studied by Liu et al. (2002); Tanha et al. (2017).
We further partition $\mathcal{T}^{\prime}_{1},\cdots,\mathcal{T}^{\prime}_{m}$ for better predictions based on $\tilde{S}\cup S_{l}$. We also repeat the following procedure recursively for each leaf $\mathcal{C}$ of $\mathcal{T}^{\prime}_{1}$:
  • Select a $\tau$-subset $\mathcal{S}$ from $d$ available features randomly without replacement;
  • Solve the optimal splitting feature $j^{*}\in\mathcal{S}$ and position $a^{*}$ w.r.t. Gini impurity and instances in $\tilde{S}\cup S_{l}$;
  • Split the current node into left and right children via the optimal splitting feature $j^{*}$ and position $a^{*}$.
The above procedure is stopped if all instances have the same label in a leaf, and get the final random tree $\mathcal{T}_{1}$. Similarly, we repeat the above procedure for $\mathcal{T}^{\prime}_{2},\cdots,\mathcal{T}^{\prime}_{m}$.
Given $m$ random trees $\mathcal{T}_{1},\cdots,\mathcal{T}_{m}$, we predict the final label of instance $\bm{x}\in\mathcal{X}$ as
$\mathcal{LF}(\bm{x})=\underset{k\in[\kappa+1]}{\arg\max}\sum_{i\in[m]}\mathcal{T}_{i,k}(\bm{x})\ ,$
where $\mathcal{T}_{i,k}(\bm{x})$ is the probability of the $k$-th category of tree $\mathcal{T}_{i}$ with leaves $\mathcal{C}_{i,1},\mathcal{C}_{i,2},\cdots,\mathcal{C}_{i,t_{i}}$, i.e.,
$\mathcal{T}_{i,k}(\bm{x})=\sum_{j\in[t_{i}]}\frac{\mathbb{I}[\bm{x}\in\mathcal{C}_{i,j}]\sum_{(\bm{x}^{\prime},y)\in S_{l}\cup\tilde{S}}\mathbb{I}[\bm{x}^{\prime}\in\mathcal{C}_{i,j},y=k]}{\max(1,\sum_{(\bm{x}^{\prime},y)\in S_{l}\cup\tilde{S}}\mathbb{I}[\bm{x}^{\prime}\in\mathcal{C}_{i,j}])}\ .$
Algorithm 1 gives detailed descriptions of our LACForest approach. Notice that our method requires the proportion $\theta$ of augmented class, which is usually unknown in practice. One feasible solution is to consider the kernel mean embedding method from Ramaswamy et al. (2016), and this is similar to previous studies on learning with augmented class Zhang et al. (2020); Shu et al. (2023).
4 Deep Neural LACForest
This section explores deep neural forests via augmented Gini impurity for learning with augmented class, which presents powerful representations with an end-to-end training manner for complex data such as images and texts. For neural forests, another advantage is to take soft splits rather than hard splits of decision trees, which could yield smoother decision boundaries and avoid overfitting Kontschieder et al. (2015). In this work ,we take one step on learning with augmented class via deep neural forests, while previous works mostly focused on all known classes Kontschieder et al. (2015); Tanno et al. (2019); Ji et al. (2020); Li et al. (2024).
4.1 Augmented Gini Impurity for Neural Trees
A deep neural forest is constructed with a DNN encoder $h$ and $m$ differentiable neural trees $\mathcal{D}\mathcal{T}_{1},\mathcal{D}\mathcal{T}_{2},\cdots,\mathcal{D}\mathcal{T}_{m}$. The key point is how to learn differentiable neural trees by incorporating information of augmented class, and our idea is to consider augmented Gini impurity as an optimization objective for neural trees.
We take neural tree $\mathcal{D}\mathcal{T}_{1}$ as an example, and it is essentially an $l$-layer complete binary tree with $l\geq 2$. We denote by $\{\mathcal{B}_{1},\cdots,\mathcal{B}_{t}\}$ and $\{\mathcal{B}_{t+1},\cdots,\mathcal{B}_{2t+2}\}$ all internal and leaf nodes of $\mathcal{D}\mathcal{T}_{1}$, respectively, where $t=2^{l-1}-1$.
For every internal node $\mathcal{B}_{i}$ ($i\in[t]$), we associate with a function $f_{i}\colon\mathcal{X}\rightarrow[0,1]$ to represent the probability of being assigned to the left child, and hence the probability of the right child is $1-f_{i}(\bm{x})$. Given encoder $h$, we define
$f_{i}(\bm{x})=\text{sigmoid}(\bm{w}_{i}^{T}h(\bm{x})+b_{i})\ ,$
where sigmoid$(z)=(1+\exp(-z))^{-1}$ and $\bm{w}_{i},b_{i}$ are learned parameters, as in Kontschieder et al. (2015).
Algorithm 2 Deep Neural LACForest
Input: Labeled data $S_{l}$, unlabeled data $S_{u}$, proportion $\theta$, hyper-parameters $m,l,T,\lambda_{\text{ce}}$
1:Randomly initialize $h$ and $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$.
2:for $j\in\{1,\cdots,T\}$ do
3:     Break $S_{l}$ and $S_{u}$ into random mini-batches.
4:     for mini-batches $S^{\prime}_{l}$ and $S^{\prime}_{u}$ from $S_{l}$ and $S_{u}$ do
5:         Compute $\mathcal{L}$ by Eqn. (11) and compute gradients.
6:         Update $h$ and $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$ by SGD.
7:     end for
8:end for
9:return $\mathcal{DF}(\bm{x})=\underset{k\in[\kappa+1]}{\arg\max}\sum_{i\in[m]}\mathcal{D}\mathcal{T}_{i,k}(\bm{x})$.
For every leaf node $\mathcal{B}_{i}$ $(i\in[2t+2]\setminus[t])$, let $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$ be the event that $\bm{x}$ is assigned to leaf $\mathcal{B}_{i}$, and denote by
$\mu_{\mathcal{B}_{i}}(\bm{x})$ $=$ $\Pr\left[\mathcal{I}(\bm{x}\to\mathcal{B}_{i})\right]$
$=$ $\prod_{j\in[t]}f_{j}(\bm{x})^{\mathbb{I}[\mathcal{B}_{i}\swarrow\mathcal{B}_{j}]}(1-f_{j}(\bm{x}))^{\mathbb{I}[\mathcal{B}_{j}\searrow\mathcal{B}_{i}]}\ ,$
where $\mathcal{B}_{i}\swarrow\mathcal{B}_{j}$ and $\mathcal{B}_{j}\searrow\mathcal{B}_{i}$ represent that $\mathcal{B}_{i}$ belongs to the left and right subtrees of $\mathcal{B}_{j}$, respectively. We also denote by
$n_{\mathcal{B}_{i},u}=\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\quad\text{and}\quad n_{\mathcal{B}_{i},l}=\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\ ,$
for unlabeled data $S_{u}$ and labeled data $S_{l}$, respectively.
Definition 2.
For $i\in[2t+2]\setminus[t]$, we define the augmented Gini impurity w.r.t. $S_{l}$ and $S_{u}$ as
$\mathcal{G}_{\mathcal{B}_{i}}(S_{l},S_{u})=1-\sum_{k\in[\kappa+1]}\vartheta^{2}_{\mathcal{B}_{i},k}(S_{l},S_{u})\ ,$ (8)
where $\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})$ is defined as
$\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})=\left(1-\frac{(1-\theta)n_{u}n_{\mathcal{B}_{i},l}}{n_{l}n_{\mathcal{B}_{i},u}}\right)_{+}\ ,$
and for $k\in[\kappa]$, $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})$ is given by
$(1-\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u}))\sum_{(\bm{x},y)\in S_{l}}\frac{\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{\mathcal{B}_{i},l}}\mathbb{I}[y=k]\ .$
Datasets #inst #feat #class Datasets #inst #feat #class Datasets #inst #feat #class Datasets #inst #feat #class
segment 2,310 19 7 mfcc 7,195 22 4 drybean 13,661 16 7 mnist 70,000 784 10
texture 5,500 40 11 usps 9,298 256 10 letter 20,000 16 26 fmnist 70,000 784 10
optdigits 5,620 64 10 har 10,299 562 6 shuttle 58,000 9 7 kuzushiji 70,000 784 10
satimage 6,435 36 6 mapping 10,845 28 6 drive 58,509 48 11 svhn 99,289 3072 10
landset 6,435 73 6 pendigits 10,992 16 10 senseveh 61,581 100 3 cifar10 60,000 3072 10
Table 1: Details of datasets.
This definition follows the augmented Gini impurity of decision trees in Definition 1, whereas deep neural trees consider probabilistic assignment of instances to nodes while the latter focuses on deterministic assignment. Essentially, $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})$ estimates the probability of the $k$-th class under the condition $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$. Based on Assumption 1, we have
Lemma 2.
For $k\in[\kappa+1]$, $\delta\in(0,1)$ and $i\in[2t+2]\setminus[t]$, we have, with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$,
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]-\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})\right|$
$\leq$ $O\left(\frac{\sqrt{\ln(1/\delta)}}{\gamma^{2}/(\gamma+1)}\left(\frac{1}{\sqrt{n_{l}}}+\frac{1}{\sqrt{n_{u}}}\right)\right)\ ,$
if ${\mathbb{E}}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\geq\gamma$, $\mathbb{E}_{\mathcal{D}_{\text{kc}}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\geq\gamma$, $n_{\mathcal{B}_{i},u}\geq\gamma n_{u}$ and $n_{\mathcal{B}_{i},l}\geq\gamma n_{l}$ for some constant $\gamma\in(0,1)$.
Denote by $\mathcal{L}^{*}_{\mathcal{B}_{i}}$ the optimal squared loss under $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$:
$\mathcal{L}^{*}_{\mathcal{B}_{i}}=\min_{\bm{w}\in\Delta_{\kappa+1}}E_{(\bm{x},y)\sim\mathcal{D}}\left[\|\bm{w}-\tilde{\bm{y}}\|^{2}_{2}|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})\right]\ ,$ (9)
where $\tilde{\bm{y}}\in\mathbb{R}^{\kappa+1}$ is the one-hot encoding of $y\in[\kappa+1]$. Based on Assumption 1 and Lemma 2, we have
Theorem 2.
For $\delta\in(0,1)$ and $i\in[2t+2]\setminus[t]$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
$\left|\mathcal{L}^{*}_{\mathcal{B}_{i}}-\mathcal{G}_{\mathcal{B}_{i}}(S_{l},S_{u})\right|\leq O\left(\frac{\kappa\sqrt{\ln(\kappa/\delta)}}{\gamma^{2}/(\gamma+1)}\left(\frac{1}{\sqrt{n_{l}}}+\frac{1}{\sqrt{n_{u}}}\right)\right)\ ,$
if ${\mathbb{E}}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\geq\gamma$, $\mathbb{E}_{\mathcal{D}_{\text{kc}}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\geq\gamma$, $n_{\mathcal{B}_{i},u}\geq\gamma n_{u}$ and $n_{\mathcal{B}_{i},l}\geq\gamma n_{l}$ for some constant $\gamma\in(0,1)$.
This theorem shows that $\mathcal{G}_{\mathcal{B}_{i}}(S_{l},S_{u})$ could be seen as an unbiased estimation of $\mathcal{L}^{*}_{\mathcal{B}_{i}}$. The detailed proof is presented in Appendix A.4, which is similar to that of Theorem 1, but takes different analysis on the convergence of $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})$ by considering the probabilistic assignment of neural trees.
4.2 Deep Neural LACForest
In the training phase, we randomly initialize encoder $h$ and neural trees $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$, and update the encoder and neural trees iteratively. In each iteration, we receive two mini-batches of labeled data $S^{\prime}_{l}$ and unlabeled data $S^{\prime}_{u}$ sampled randomly from $S_{l}$ and $S_{u}$, respectively. We introduce
$\mathcal{L}_{\text{ag}}(\mathcal{D}\mathcal{T}_{i})=\sum_{j\in[2t+2]\setminus[t]}\omega_{i,j}(S^{\prime}_{u})\mathcal{G}_{\mathcal{B}_{i,j}}(S^{\prime}_{l},S^{\prime}_{u})\ ,$ (10)
where $\{\mathcal{B}_{i,j}\}_{j\in[2t+2]\setminus[t]}$ denote all leaf nodes of $\mathcal{D}\mathcal{T}_{i}$, and $\omega_{i,j}(S^{\prime}_{u})$ is the proportion of unlabeled data in $\mathcal{B}_{i,j}$, i.e.,
$\omega_{i,j}(S^{\prime}_{u})=\frac{1}{|S^{\prime}_{u}|}\sum_{\bm{x}\in S^{\prime}_{u}}{\mu_{\mathcal{B}_{i,j}}(\bm{x})}\ .$
The augmented Gini loss in Eqn. (10) is a weighted average augmented Gini impurity for all leaf nodes of $\mathcal{D}\mathcal{T}_{i}$.
For better representation, we introduce an auxiliary linear predictor $g\colon\mathbb{R}^{d^{\prime}}\rightarrow\Delta_{\kappa}$ as done in Yan et al. (2021), and take the cross entropy loss on known classes w.r.t. $S^{\prime}_{l}$ as
$\mathcal{L}_{\text{ce}}=-\frac{1}{|S^{\prime}_{l}|}\sum_{(\bm{x},y)\in S^{\prime}_{l}}\sum_{t\in[\kappa]}\tilde{y}_{t}\log\left(g_{t}\left(h\left(\bm{x}\right)\right)\right)\ .$
We finally get the optimization objective as follows:
$\mathcal{L}=\sum_{i\in[m]}\frac{\mathcal{L}_{\text{ag}}(\mathcal{D}\mathcal{T}_{i})}{m}+\lambda_{\text{ce}}\mathcal{L}_{\text{ce}}\ ,$ (11)
where $\lambda_{\text{ce}}>0$ is a trade-off parameter. Intuitively, the cross entropy loss is beneficial to learning a basic representation of target data, and the augmented gini loss could improve the ability of neural trees to distinguish augmented class from known classes, as shown by Theorem 2.
Datasets Our LACForest GLAC EULAC EVM PAC-iForest OSNN LACU-SVM OVR-SVM
segment .9436$\pm$.0186 .8838$\pm$.0390$\bullet$ .9256$\pm$.0276$\bullet$ .8612$\pm$.0746$\bullet$ .6678$\pm$.0763$\bullet$ .5467$\pm$.0384$\bullet$ .4918$\pm$.0796$\bullet$ .6109$\pm$.0901$\bullet$
texture .9151$\pm$.0176 .9130$\pm$.0328 .9138$\pm$.0208 .8724$\pm$.0270$\bullet$ .7107$\pm$.0518$\bullet$ .5956$\pm$.0405$\bullet$ .5829$\pm$.0478$\bullet$ .6078$\pm$.0587$\bullet$
optdigits .9260$\pm$.0203 .8842$\pm$.0384$\bullet$ .9243$\pm$.0186 .9069$\pm$.0234$\bullet$ .7202$\pm$.0415$\bullet$ .6572$\pm$.0282$\bullet$ .7385$\pm$.0326$\bullet$ .7547$\pm$.0552$\bullet$
satimage .8791$\pm$.0322 .8215$\pm$.0526$\bullet$ .8644$\pm$.0371$\bullet$ .7238$\pm$.0664$\bullet$ .7460$\pm$.0566$\bullet$ .5101$\pm$.0380$\bullet$ .6324$\pm$.0165$\bullet$ .5228$\pm$.0568$\bullet$
landset .9243$\pm$.0206 .8627$\pm$.0262$\bullet$ .8857$\pm$.0236$\bullet$ .8120$\pm$.0464$\bullet$ .7572$\pm$.0671$\bullet$ .5302$\pm$.0257$\bullet$ .6289$\pm$.0124$\bullet$ .5366$\pm$.0316$\bullet$
mfcc .9418$\pm$.0144 .8937$\pm$.0194$\bullet$ .9506$\pm$.0137$\circ$ .8669$\pm$.0888$\bullet$ .7901$\pm$.0710$\bullet$ .5329$\pm$.0311$\bullet$ .7751$\pm$.0362$\bullet$ .6428$\pm$.0215$\bullet$
usps .8931$\pm$.0198 .8645$\pm$.0382$\bullet$ .8955$\pm$.0250 .7959$\pm$.0697$\bullet$ .5695$\pm$.0894$\bullet$ .6413$\pm$.0317$\bullet$ .7642$\pm$.0379$\bullet$ .7488$\pm$.0464$\bullet$
har .9020$\pm$.0319 .8772$\pm$.0460$\bullet$ .8922$\pm$.0269$\bullet$ .5016$\pm$.0558$\bullet$ .5963$\pm$.0699$\bullet$ .5048$\pm$.0386$\bullet$ .5570$\pm$.0408$\bullet$ .5046$\pm$.0251$\bullet$
mapping .8612$\pm$.0199 .7964$\pm$.0350$\bullet$ .8515$\pm$.0309$\bullet$ .7111$\pm$.1320$\bullet$ .6712$\pm$.1046$\bullet$ .6218$\pm$.0605$\bullet$ .6538$\pm$.0559$\bullet$ .5210$\pm$.0596$\bullet$
pendigits .9281$\pm$.0215 .8872$\pm$.0241$\bullet$ .9276$\pm$.0245 .9266$\pm$.0330 .7709$\pm$.0742$\bullet$ .5822$\pm$.0243$\bullet$ .7352$\pm$.0513$\bullet$ .6424$\pm$.0469$\bullet$
drybean .8932$\pm$.0208 .9026$\pm$.0528$\circ$ .9038$\pm$.0202$\circ$ .7834$\pm$.0428$\bullet$ .7555$\pm$.0693$\bullet$ .5661$\pm$.0580$\bullet$ .6384$\pm$.0397$\bullet$ .5471$\pm$.0478$\bullet$
letter .7402$\pm$.0335 .6331$\pm$.0394$\bullet$ .6057$\pm$.0387$\bullet$ .7004$\pm$.0460$\bullet$ .4875$\pm$.0373$\bullet$ .5980$\pm$.0231$\bullet$ .5547$\pm$.0538$\bullet$ .6238$\pm$.0399$\bullet$
shuttle .9750$\pm$.0197 .9444$\pm$.0506$\bullet$ .9770$\pm$.0116 .6245$\pm$.0347$\bullet$ .6042$\pm$.0326$\bullet$ .5464$\pm$.0649$\bullet$ .6734$\pm$.0218$\bullet$ .4862$\pm$.0249$\bullet$
drive .8529$\pm$.0532 .6114$\pm$.0824$\bullet$ .7844$\pm$.0460$\bullet$ .7708$\pm$.0476$\bullet$ .4296$\pm$.0578$\bullet$ .5474$\pm$.0246$\bullet$ .6032$\pm$.0783$\bullet$ .5287$\pm$.1040$\bullet$
senseveh .7977$\pm$.0295 .7545$\pm$.0406$\bullet$ .7925$\pm$.0196$\bullet$ .5685$\pm$.0572$\bullet$ .5580$\pm$.1089$\bullet$ .5331$\pm$.0499$\bullet$ .5973$\pm$.0425$\bullet$ .5726$\pm$.0688$\bullet$
mnist .8540$\pm$.0299 .7947$\pm$.0408$\bullet$ .8390$\pm$.0380$\bullet$ .6181$\pm$.0475$\bullet$ .5369$\pm$.0827$\bullet$ .6504$\pm$.0386$\bullet$ .6481$\pm$.0496$\bullet$ .7234$\pm$.0505$\bullet$
fmnist .8008$\pm$.0269 .7592$\pm$.0343$\bullet$ .7672$\pm$.0382$\bullet$ .6508$\pm$.0368$\bullet$ .6090$\pm$.0577$\bullet$ .5152$\pm$.0534$\bullet$ .5674$\pm$.0629$\bullet$ .5499$\pm$.0727$\bullet$
average .8840$\pm$.0591 .8285$\pm$.0916 .8648$\pm$.0862 .7468$\pm$.1204 .6459$\pm$.1042 .5694$\pm$.0486 .6378$\pm$.0774 .5955$\pm$.0822
win/tie/loss 15/1/1 10/5/2 16/1/0 17/0/0 17/0/0 17/0/0 17/0/0
Table 2: Experimental comparisons of accuracy (mean±$\pm$±std) for compared methods, and $\bullet$/$\circ$ indicates that our approach is significantly better/worse than the corresponding method (paired $t$-test at $95\%$ significance level).
Notice that the optimization objective $\mathcal{L}$ in Eqn. (11) is differentiable w.r.t. the parameters of feature encoder $h$ and neural trees $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$, and we could update the entire model with stochastic gradient descent directly.
Given feature encoder $h$ and neural trees $\mathcal{D}\mathcal{T}_{1},\cdots\mathcal{D}\mathcal{T}_{m}$, we could present the prediction for instance $\bm{x}\in\mathcal{X}$ as
$\mathcal{DF}(\bm{x})=\underset{k\in[\kappa+1]}{\arg\max}\sum_{i\in[m]}\mathcal{D}\mathcal{T}_{i,k}(\bm{x})\ ,$
where $\mathcal{D}\mathcal{T}_{i,k}(\bm{x})$ shows the probability of the $k$-th class w.r.t. neural tree $\mathcal{D}\mathcal{T}_{i}$, i.e.,
$\mathcal{D}\mathcal{T}_{i,k}(\bm{x})=\sum_{j\in[2t]\setminus[t-1]}\mu_{\mathcal{B}_{i,j}}(\bm{x})\vartheta_{\mathcal{B}_{i,j},k}(S_{l},S_{u})\ ,$
with $\vartheta_{\mathcal{B}_{i,j},k}(S_{l},S_{u})$ given by Definition 2.
Algorithm 2 presents the detailed description of our deep neural LACForest approach. In experiments, we should take relatively large batch sizes for mini-batches $S^{\prime}_{l}$ and $S^{\prime}_{u}$, since $\mathcal{G}_{\mathcal{B}_{i,j}}(S^{\prime}_{l},S^{\prime}_{u})$ in Eqn. (10) converges to the optimal squared loss in the rate of $O(1/\sqrt{\min\{|S^{\prime}_{l}|,|S^{\prime}_{u}|\}})$ from Theorem 2.
4.3 Related Works
Zhou and Chen Zhou and Chen (2002) introduced the problem of class-incremental learning via a few labeled augmented instances. Fink et al. Fink et al. (2006) learned multiple binary classifiers for this problem and Topalis and Polikar Topalis and Polikar (2008) considered voting classifiers. Recent years have witnessed increasing attention on the design of practical algorithms for this problem Li and Hoiem (2017); Rebuffi et al. (2017); Yan et al. (2021); Zou et al. (2022); Zhou et al. (2024). These methods can not be applied to our learning scenario directly, since we have no access to any labeled augmented instances.
Another relevant problem is open-set recognition in computer vision. Scheirer et al. Scheirer et al. (2013) introduced open space risk to penalize predictions outside the support of training data. Along this line, various approaches have been developed based on open space risk Scheirer et al. (2014), extreme value theory Bendale and Boult (2016); Rudd et al. (2017), nearest neighbors Mendes Júnior et al. (2017) and generative neural networks Ge et al. (2017); Neal et al. (2018); Chen et al. (2021). Those studies are strongly based on some geometric assumptions.
Figure 2: Experimental comparisons of AUC on the detection of augmented class.
5 Experiments
We conduct experiments on 15 benchmark datasets and 5 image datasets, and the details are summarized in Table 1. Most datasets have been well-studied in previous works on learning with augmented class.
5.1 Evaluation of Our LACForest Approach
For our LACForest approach, we compare with seven state-of-the-art approaches: GLAC Shu et al. (2023), EULAC Zhang et al. (2020), PAC-iForest Liu et al. (2018), EVM Rudd et al. (2017), OSNN Mendes Júnior et al. (2017), LACU-SVM Da et al. (2014) and OVR-SVM Rifkin and Klautau (2004). More details on these approaches could be found in Appendix B.1.
For each dataset, we randomly select half of classes as the augmented class with the rest as known classes, following Zhang et al. (2020). We then randomly sample 500 examples of known classes as labeled data $S_{l}$, and 1000 instances as unlabeled data $S_{u}$ and 100 instances as testing data. We take $\theta=0.5$ in Eqn. (1), and more experimental settings could be found in Appendix B.1. The performance is evaluated by 10 trials of random selections of augmented class, and with 10 times of random data sampling. The average test accuracies are obtained over these 100 runs, as shown in Table 2.
Datasets Our approach Deep-GLAC Deep-EULAC ARPL G-Openmax OSRCI Openmax Softmax-T
mnist .9844$\pm$.0021 .9778$\pm$.0039 .9596$\pm$.0033 .9304$\pm$.0203 .8934$\pm$.0064 .9114$\pm$.0047 .8876$\pm$.0042 .8834$\pm$.0029
fmnist .9024$\pm$.0114 .9010$\pm$.0148 .8464$\pm$.0085 .7682$\pm$.0102 .6820$\pm$.0148 .6912$\pm$.0070 .6672$\pm$.0139 .5878$\pm$.0054
kuzushiji .9636$\pm$.0044 .9516$\pm$.0032 .8872$\pm$.0032 .9002$\pm$.0143 .8570$\pm$.0055 .8602$\pm$.0046 .8426$\pm$.0070 .8282$\pm$.0069
svhn .9238$\pm$.0134 .8926$\pm$.0168 .8330$\pm$.0087 .8060$\pm$.0069 .7868$\pm$.0161 .7912$\pm$.0072 .7888$\pm$.0057 .7252$\pm$.0047
cifar10 .8008$\pm$.0331 .7840$\pm$.0426 .7168$\pm$.0231 .7208$\pm$.0032 .6560$\pm$.0125 .6788$\pm$.0084 .6612$\pm$.0293 .6350$\pm$.0203
average .9150$\pm$.0639 .9016$\pm$.0669 .8486$\pm$.0793 .8251$\pm$.0790 .7750$\pm$.0933 .7866$\pm$.0914 .7695$\pm$.0915 .7319$\pm$.1117
Table 3: Experimental comparisons of accuracy (mean±$\pm$±std) over 5 image datasets, and the best performance is highlighted in bold.
Figure 3: Evaluations of LACForest over different proportions of augmented data. The larger the curve, the better the performance.
Figure 4: Evaluations of LACForest under different sizes of unlabeled data $S_{u}$. The larger the curve, the better the performance.
It is clear that our LACForest method achieves significantly better performance than previous EVM, OSNN, PAC-iForest and OVR-SVM, as our LACForest wins in most times and never loses. This is because those methods mainly focus on labeled data from known classes, but without exploring information from unlabeled data.
Our LACForest also outperforms LACU-SVM, which is heavily dependent on the low-separation assumption over data. In comparison to GLAC and EULAC, our LACForest achieves better and comparable performance in most times, except for datasets mfcc and drybean, partially because of class imbalance in these datasets, which makes it difficult to accurately estimate proportions of augmented class in nodes.
We also take the average AUC to show the performance on the detection of augmented class in Figure 2. It is obvious that our LACForest takes better and comparable performance on the detection of augmented class over most datasets, since our method could explore augmented class effectively in each leaf node, as shown by Lemma 1.
Figure 5: Influence of parameter $\lambda_{\text{ce}}$.
5.2 Evaluation of Our Deep Neural LACForest
For our deep neural LACForest, we compare with seven deep learning methods for augmented class: Deep-GLAC Shu et al. (2023), ARPL Chen et al. (2021), Deep-EULAC Zhang et al. (2020), OSRCI Neal et al. (2018), G-Openmax Ge et al. (2017), Openmax Bendale and Boult (2016) and Softmax-T Hendrycks and Gimpel (2016). More details on these approaches are presented in Appendix B.2.
We take a three-layer convolutional neural network as the backbone neural network on mnist, fmnist and kuzushiji, and consider VGG16 Simonyan and Zisserman (2015) on svhn and cifar10, as done in Shu et al. (2023). We randomly select four classes as the augmented class and take the rest as known classes, and set $\theta=0.4$ similarly to Shu et al. (2023).
We take average test accuracies over 5 random selections of augmented class as our performance measure, and the experimental comparisons are shown in Table 3. It is clear that our approach takes better performance than ARPL, G-Openmax, OSRCI, Openmax and Softmax-T, due to some additional geometric assumptions over those methods. For Deep-GLAC and Deep-EULAC, our method achieves better and comparable performance in most times, and an intuitive explanation is that our approach could effectively explore augmented class from each local region during tree partitions.
Parameter Influence. We analyze the influence of various parameters on several datasets, and the trends are similar for other datasets. Figure 3 shows that our LACForest gets better performance under different proportions $\theta\in[0.3,0.8]$ for augmented data. Figure 4 shows the performance with different sizes of unlabeled data, where our LACForest achieves better and stable performance with the increase of unlabeled data, in consistency with Theorem 1. For deep neural LACForest, Figure 5 shows that our approach is insensitive to parameter $\lambda_{\text{ce}}$ and generally works well for $\lambda_{\text{ce}}\in[0.2,2]$. Figure 6 shows the influence of the depth of neural trees, and our method takes stable results when the tree depth $l\geq 5$.
Figure 6: Influence of neural trees’ depth.
6 Conclusion
This work studies learning with augmented class via forests, where an augmented class may appear in testing data yet not in training data. We introduce the augmented Gini impurity as a new splitting criterion by incorporating information of augmented class during tree construction. We develop the approach on Learning with Augmented Class via Forests, which constructs shallow forests based on augmented Gini impurity and further splits forests with pseudo-labeled augmented instances. We also explore deep neural forests with a new optimization objective via our augmented Gini impurity. We validate the effectiveness of our methods both empirically and theoretically. An interesting future direction is to exploit our methods under other learning settings such as streaming datasets and multiple augmented classes.
Acknowledgments
The authors want to thank the reviewers for their helpful comments and suggestions. This research was supported by National Key R&D Program of China (2021ZD0112802) and NSFC (62376119). Wei Gao is the corresponding author.
Appendix A Detailed Theoretical Proofs
A.1 Proof of Lemma 1
Proof.
Recall that $S_{\mathcal{C},l}=\{(\bm{x},y)\colon(\bm{x},y)\in S_{l}\text{ and }\bm{x}\in\mathcal{C}\}$, $S_{\mathcal{C},u}=S_{u}\cap\mathcal{C}$, $n_{\mathcal{C},l}=|S_{\mathcal{C},l}|$ and $n_{\mathcal{C},u}=|S_{\mathcal{C},u}|$. According to the Hoeffding’s inequality, we have, for some $\delta_{1}\in(0,\frac{1}{3})$
$\left|\frac{n_{C,l}}{n_{l}}-\Pr_{(\bm{x},y)\sim\mathcal{D}_{\text{kc}}}[\bm{x}\in C]\right|\leq\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}\ ,$ (12)
with probability at least $1-\delta_{1}$. For simplicity, we introduce $n^{\prime}_{\mathcal{C},u}$ as follows:
$n^{\prime}_{\mathcal{C},u}=\sum_{\bm{x}\in S_{u}}\mathbb{I}[\bm{x}\in\mathcal{C}]\Pr_{(\bm{x}^{\prime},y^{\prime})\sim\mathcal{D}}\left[y^{\prime}\in[\kappa]|\bm{x}^{\prime}=\bm{x}\right]\ .$ (13)
Let $f_{\mathcal{X}}(\bm{x})$ and $f_{\mathcal{X},\mathcal{Y}}(\bm{x},y)$ be the probability density functions of distribution $\mathcal{D}_{\mathcal{X}}$ and $\mathcal{D}$, respectively, and we have
$\underset{\bm{x}\sim\mathcal{D}_{\mathcal{X}}}{E}\left[\mathbb{I}[\bm{x}\in\mathcal{C}]\Pr_{(\bm{x}^{\prime},y^{\prime})\sim\mathcal{D}}\left[y^{\prime}\in[\kappa]|\bm{x}^{\prime}=\bm{x}\right]\right]$
[…]
$=\leavevmode\nobreak\ \int_{\bm{x}\sim\mathcal{D}_{\mathcal{X}}}\mathbb{I}[\bm{x}\in\mathcal{C}]\sum_{k\in[\kappa]}f_{\mathcal{X},\mathcal{Y}}(\bm{x},k)d\bm{x}\ ,$
and this follows that, from that $\mathcal{D}_{\mathcal{X}}$ is the marginal distribution of $\mathcal{D}$ over $\mathcal{X}$
$\underset{\bm{x}\sim\mathcal{D}_{\mathcal{X}}}{E}\left[\mathbb{I}[\bm{x}\in\mathcal{C}]\Pr_{(\bm{x}^{\prime},y^{\prime})\sim\mathcal{D}}\left[y^{\prime}\in[\kappa]|\bm{x}^{\prime}=\bm{x}\right]\right]$
[…]
$=\leavevmode\nobreak\ \Pr_{(\bm{x},y)\sim\mathcal{D}}\left[y\in[\kappa],\bm{x}\in\mathcal{C}\right]\ .$ (14)
From Eqns. (13)-(14) and the Hoeffding’s inequality, we have, for some $\delta_{2}\in(0,\frac{1}{3})$
$\left|\frac{n^{\prime}_{\mathcal{C},u}}{n_{u}}-\underset{\bm{x}\sim\mathcal{D}_{\mathcal{X}}}{E}\left[\mathbb{I}[\bm{x}\in\mathcal{C}]\Pr_{(\bm{x}^{\prime},y^{\prime})\sim\mathcal{D}}\left[y^{\prime}\in[\kappa]|\bm{x}^{\prime}=\bm{x}\right]\right]\right|$
$=\leavevmode\nobreak\ \left|\frac{n^{\prime}_{\mathcal{C},u}}{n_{u}}-\Pr_{(\bm{x},y)\sim\mathcal{D}}\left[y\in[\kappa],\bm{x}\in\mathcal{C}\right]\right|\leq\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}}\ ,$ (15)
with probability at least $1-\delta_{2}$. From Eqn. (1), we have
$\Pr_{(\bm{x},y)\sim\mathcal{D}}\left[y\in[\kappa],\bm{x}\in\mathcal{C}\right]=\Pr_{(\bm{x},y)\sim\mathcal{D}}\left[y\in\left[\kappa\right]\right]\Pr_{(\bm{x},y)\sim\mathcal{D}}\left[\bm{x}\in\mathcal{C}|y\in[\kappa]\right]=(1-\theta)\Pr_{(\bm{x},y)\sim\mathcal{D}_{\text{kc}}}\left[\bm{x}\in\mathcal{C}\right]\ ,$
and it follows that, from Eqn. (15)
$\left|\frac{n^{\prime}_{\mathcal{C},u}}{(1-\theta)n_{u}}-\Pr_{(\bm{x},y)\sim\mathcal{D}_{\text{kc}}}[\bm{x}\in\mathcal{C}]\right|\leq\frac{1}{1-\theta}\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}}\ ,$ (16)
with probability at least $1-\delta_{2}$. Combining Eqns. (12) and (16), we have
$\left|\frac{n_{\mathcal{C},l}}{n_{l}}-\frac{n^{\prime}_{\mathcal{C},u}}{(1-\theta)n_{u}}\right|\leq\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}+\frac{1}{1-\theta}\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}}\ ,$ (17)
with probability at least $1-\delta_{1}-\delta_{2}$. From Eqn. (13) and the Hoeffding’s inequality, we have
$\left|\left(1-\frac{n^{\prime}_{\mathcal{C},u}}{n_{\mathcal{C},u}}\right)-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\bm{x}\in\mathcal{C}]\right|$
$=\leavevmode\nobreak\ \left|\frac{n^{\prime}_{\mathcal{C},u}}{n_{\mathcal{C},u}}-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y\in[\kappa]|\bm{x}\in\mathcal{C}]\right|\leq\sqrt{\frac{\ln(2/\delta_{3})}{2n_{\mathcal{C},u}}}\ ,$ (18)
with probability at least $1-\delta_{3}$ with some $\delta_{3}\in(0,\frac{1}{3})$. From $n_{\mathcal{C},u}\geq\gamma n_{u}>0$ and Definition 1, we have
$\vartheta_{C,\kappa+1}(S_{l},S_{u})=\mathbb{I}[n_{\mathcal{C},u}>0]\left(1-\frac{\left(1-\theta\right)n_{u}n_{\mathcal{C},l}}{n_{l}n_{\mathcal{C},u}}\right)_{+}=\left(1-\frac{\left(1-\theta\right)n_{u}n_{\mathcal{C},l}}{n_{l}n_{\mathcal{C},u}}\right)_{+}\ .$
For simplicity, we denote by $\theta_{\mathcal{C}}=\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\bm{x}\in\mathcal{C}]$, and this follows that
$\left|\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})-\theta_{\mathcal{C}}\right|$
$\leq\leavevmode\nobreak\ \left|\left(1-\frac{\left(1-\theta\right)n_{u}n_{\mathcal{C},l}}{n_{l}n_{\mathcal{C},u}}\right)-\theta_{\mathcal{C}}\right|$
$\leq\leavevmode\nobreak\ \left|\left(1-\frac{\left(1-\theta\right)n_{u}n_{\mathcal{C},l}}{n_{l}n_{\mathcal{C},u}}\right)-\left(1-\frac{n^{\prime}_{\mathcal{C},u}}{n_{\mathcal{C},u}}\right)\right|+\left|\left(1-\frac{n^{\prime}_{\mathcal{C},u}}{n_{\mathcal{C},u}}\right)-\theta_{\mathcal{C}}\right|$
$\leq\leavevmode\nobreak\ \left|\frac{\left(1-\theta\right)n_{u}n_{\mathcal{C},l}}{n_{l}n_{\mathcal{C},u}}-\frac{n^{\prime}_{\mathcal{C},u}}{n_{\mathcal{C},u}}\right|+\left|\left(1-\frac{n^{\prime}_{\mathcal{C},u}}{n_{\mathcal{C},u}}\right)-\theta_{\mathcal{C}}\right|$
$\leq\leavevmode\nobreak\ \frac{(1-\theta)n_{u}}{n_{\mathcal{C},u}}\left|\frac{n_{\mathcal{C},l}}{n_{l}}-\frac{n^{\prime}_{\mathcal{C},u}}{(1-\theta)n_{u}}\right|+\left|\left(1-\frac{n^{\prime}_{\mathcal{C},u}}{n_{\mathcal{C},u}}\right)-\theta_{\mathcal{C}}\right|\ ,$
and we further have, from Eqns. (17)-(18) and $n_{\mathcal{C},u}\geq\gamma n_{u}$
$\left|\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})-\theta_{\mathcal{C}}\right|$
$\leq\leavevmode\nobreak\ \frac{\left(1-\theta\right)n_{u}}{n_{\mathcal{C},u}}(\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}+\frac{1}{1-\theta}\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}})+\sqrt{\frac{\ln(2/\delta_{3})}{2n_{\mathcal{C},u}}}$
$\leq\leavevmode\nobreak\ \frac{1}{\gamma}\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}}+\sqrt{\frac{\ln(2/\delta_{3})}{2\gamma n_{u}}}\ ,$
with probability at least $1-\delta_{1}-\delta_{2}-\delta_{3}$, which completes the proof by setting $\delta_{1}=\delta_{2}=\delta_{3}=\delta/3$. ∎
A.2 Proof of Theorem 1
We begin with some useful lemmas as follows.
Lemma 3.
For real numbers $a,b,c,d$ with $a,d\in[0,1]$, we have
$\left|ab-cd\right|\leq\left|a-c\right|+\left|b-d\right|\ .$
Proof.
From the triangle inequality, we have
$\left|ab-cd\right|\leq\left|ab-ad\right|+\left|ad-cd\right|=a\left|b-d\right|+d\left|a-c\right|\ ,$
which completes the proof from $a,d\in[0,1]$. ∎
Lemma 4.
For any instance space $\mathcal{C}\subseteq\mathcal{X}$ and $\mathcal{L}^{*}_{\mathcal{C}}$ defined by Eqn. (3), we have
$\mathcal{L}^{*}_{C}=1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}\ .$
Proof.
For any $\bm{w}\in\Delta_{\kappa+1}$, we have, by some simple calculations
$E_{(\bm{x},y)\sim\mathcal{D}}\left[\mathcal{L}(\bm{w},\tilde{\bm{y}})|\bm{x}\in\mathcal{C}\right]$
$=\leavevmode\nobreak\ \sum_{k\in[\kappa+1]}\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\left(\left(1-w_{k}\right)^{2}-w^{2}_{k}+\sum_{t\in[\kappa+1]}w^{2}_{t}\right)$
= $=\leavevmode\nobreak\ 1+\sum_{k\in[\kappa+1]}\left(w_{k}-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}\ ,$
$\geq\leavevmode\nobreak\ 1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}\ ,$
and this follows that, from $\mathcal{L}^{*}_{C}=\min\limits_{\bm{w}^{*}\in\Delta_{\kappa+1}}E_{(\bm{x},y)\sim\mathcal{D}}\left[\mathcal{L}(\bm{w},\tilde{\bm{y}})|\bm{x}\in\mathcal{C}\right]$
$\mathcal{L}^{*}_{C}\geq 1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}\ .$ (19)
Let $w^{\prime}_{k}=\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in C]$ for $k\in[\kappa+1]$, and we have
$E_{(\bm{x},y)\sim\mathcal{D}}\left[\mathcal{L}(\bm{w}^{\prime},\tilde{\bm{y}})|\bm{x}\in C\right]=1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in C]\right)^{2}\geq\mathcal{L}^{*}_{C}\ ,$
which completes the proof from Eqn. (19). ∎
Lemma 5.
For instance space $\mathcal{C}\subseteq\mathcal{X}$ and $k\in[\kappa]$, if $n_{\mathcal{C},u}/n_{u}\geq\gamma$ and $n_{\mathcal{C},l}/n_{l}\geq\gamma$ for some constant $\gamma\in(0,1)$, then we have, with probability at least $1-\delta_{1}-\delta_{2}$ for some $\delta_{1},\delta_{2}>0$ with $\delta_{1}+\delta_{2}<1$
$\left|\vartheta_{\mathcal{C},k}(S_{l},S_{u})-\Pr_{(\bm{x},y)\in\mathcal{D}}\left[y=k|\bm{x}\in\mathcal{C}\right]\right|\leq\sqrt{\frac{\ln(6/\delta_{1})}{2\gamma n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{l}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{u}}}+\sqrt{\frac{\ln(2/\delta_{2})}{2\gamma n_{l}}}\ .$
Proof.
For simplicity, we denote by $\theta_{\mathcal{C}}=\Pr_{(\bm{x},y)\in\mathcal{D}}\left[y=\kappa+1|\bm{x}\in\mathcal{C}\right]$. If $\theta_{\mathcal{C}}=1$, then we have, for $k\in[\kappa]$
$\Pr_{(\bm{x},y)\in\mathcal{D}}\left[y=k|\bm{x}\in\mathcal{C}\right]=0\ ,$ (20)
and this follows that, from Definition 1 and $\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]=0$
$\vartheta_{C,k}(S_{l},S_{u})=(1-\vartheta_{\mathcal{C},\kappa+1})\frac{\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]}{\min(n_{\mathcal{C},l},1)}=0\ ,$
and we further have, from Eqn. (20)
$\left|\vartheta_{\mathcal{C},k}(S_{l},S_{u})-\Pr_{(\bm{x},y)\in\mathcal{D}}\left[y=k|\bm{x}\in\mathcal{C}\right]\right|=0\ .$ (21)
On the other hand, if $\theta_{\mathcal{C}}\in[0,1)$, then we have, from Lemma 1
$\left|\vartheta_{C,\kappa+1}(S_{l},S_{u})-\theta_{\mathcal{C}}\right|\leq\sqrt{\frac{\ln(6/\delta_{1})}{2\gamma n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{l}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{u}}}\ ,$ (22)
with probability at least $1-\delta_{1}$. From the Hoeffding’s inequality and $n_{\mathcal{C},l}\geq\gamma n_{l}$, we also have
$\left|\frac{1}{n_{\mathcal{C},l}}\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]-\Pr_{(\bm{x},y)\sim\mathcal{D}_{\text{kc}}}[y=k|\bm{x}\in\mathcal{C}]\right|\leq\sqrt{\frac{\ln(2/\delta_{2})}{2n_{C,l}}}\leq\sqrt{\frac{\ln(2/\delta_{2})}{2\gamma n_{l}}}\ ,$ (23)
with probability at least $1-\delta_{2}$. Notice that $\mathcal{D}_{\text{kc}}$ is the marginal distribution of $\mathcal{D}$ over $\mathcal{X}\times[\kappa]$, and we have
$\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]=\Pr_{(\bm{x},y)\sim\mathcal{D}}[y\in[\kappa]|\bm{x}\in\mathcal{C}]\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|y\in[\kappa],\bm{x}\in\mathcal{C}]=(1-\theta_{\mathcal{C}})\Pr_{(\bm{x},y)\sim\mathcal{D}_{\text{kc}}}[y=k|\bm{x}\in\mathcal{C}]\ ,$
and this follows that, from Eqn. (23) and $\theta_{\mathcal{C}}\in[0,1)$
$\left|\frac{1}{n_{\mathcal{C},l}}\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]-\frac{1}{1-\theta_{C}}\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right|\leq\sqrt{\frac{\ln(2/\delta_{2})}{2\gamma n_{l}}}\ ,$ (24)
with probability at least $1-\delta_{2}$. From Definition 1 and $n_{\mathcal{C},l}\geq\gamma n_{l}\geq 1$, we have
$\vartheta_{\mathcal{C},k}(S_{l},S_{u})=\frac{1-\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})}{\min(n_{\mathcal{C},l},1)}\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]=\frac{1-\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})}{n_{\mathcal{C},l}}\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]\ ,$
and this follows that, from Lemma 3
$\left|\vartheta_{\mathcal{C},k}(S_{l},S_{u})-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right|$
$\leq\leavevmode\nobreak\ \left|\left(1-\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})\right)\frac{\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]}{n_{\mathcal{C},l}}-(1-\theta_{\mathcal{C}})\frac{\Pr\limits_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]}{1-\theta_{\mathcal{C}}}\right|$
$\leq\leavevmode\nobreak\ \left|\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})-\theta_{\mathcal{C}}\right|+\left|\frac{\sum_{(\bm{x},y)\in S_{l}}\mathbb{I}[\bm{x}\in\mathcal{C}\wedge y=k]}{n_{\mathcal{C},l}}-\frac{\Pr\limits_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]}{1-\theta_{\mathcal{C}}}\right|\ ,$
and we further have, from Eqns.(22) and (24)
$\left|\vartheta_{\mathcal{C},k}(S_{l},S_{u})-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right|\leq\sqrt{\frac{\ln(6/\delta_{1})}{2\gamma n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{l}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{u}}}+\sqrt{\frac{\ln(2/\delta_{2})}{2\gamma n_{l}}}\ .$
This completes the proof by combining with Eqn. (21). ∎
Proof of Theorem 1. From Lemma 4, we have
$\mathcal{L}^{*}_{\mathcal{C}}=1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}\ ,$
and this follows that, from Eqn. (2)
$\left|\mathcal{L}^{*}_{\mathcal{C}}-\mathcal{G}_{\mathcal{C}}(S_{l},S_{u})\right|\leq\sum_{k\in[\kappa+1]}\left|\vartheta^{2}_{\mathcal{C},k}(S_{l},S_{u})-\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}\right|\ .$ (25)
For $k=\kappa+1$, we have, from $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})\in[0,1]$ and $\theta_{\mathcal{C}}=\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\bm{x}\in\mathcal{C}]\in[0,1]$
$\left|\vartheta^{2}_{\mathcal{C},\kappa+1}(S_{l},S_{u})-\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\bm{x}\in\mathcal{C}]\right)^{2}\right|$
$\leq\leavevmode\nobreak\ \left|\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})-\theta_{\mathcal{C}}\right|\times\left|\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})+\theta_{\mathcal{C}}\right|\leq 2\left|\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})-\theta_{\mathcal{C}}\right|\ ,$
and it follows that, from Lemma 1
$\left|\vartheta^{2}_{\mathcal{C},\kappa+1}(S_{l},S_{u})-\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\bm{x}\in\mathcal{C}]\right)^{2}\right|\leq 2\sqrt{\frac{\ln(6/\delta_{1})}{2\gamma n_{u}}}+\frac{2}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{l}}}+\frac{2}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{u}}}\ ,$ (26)
with probability at least $1-\delta_{1}$ for some $\delta_{1}\in(0,1)$. For $k\in[\kappa]$, we similarly have
$\left|\vartheta^{2}_{\mathcal{C},k}(S_{l},S_{u})-\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right)^{2}\right|\leq 2\left|\vartheta_{\mathcal{C},k}(S_{l},S_{u})-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\bm{x}\in\mathcal{C}]\right|\ ,$
and this follows that, from Eqns. (25)-(26) and Lemma 5
$\left|\mathcal{L}^{*}_{\mathcal{C}}-\mathcal{G}_{\mathcal{C}}(S_{l},S_{u})\right|\leq(2\kappa+2)\sqrt{\frac{\ln(6/\delta_{1})}{2\gamma n_{u}}}+\frac{2\kappa+2}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{l}}}+\frac{2\kappa+2}{\gamma}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{u}}}+2\kappa\sqrt{\frac{\ln(2/\delta_{2})}{2\gamma n_{l}}}\$
with probability at least $1-(\kappa+1)\delta_{1}-\kappa\delta_{2}$. We finally complete the proof by setting $\delta_{1}=\frac{\delta}{2(\kappa+1)}$ and $\delta_{2}=\frac{\delta}{2\kappa}$ and some simple algebraic calculations. ∎
A.3 Proof of Lemma 2
We first introduce a necessary Lemma as follows.
Lemma 6.
For $i\in[2t+2]\setminus[t]$ and $\delta\in(0,1)$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})\right|\leq\frac{1-\theta}{\gamma}\sqrt{\frac{\ln(8/\delta)}{2n_{l}}}+\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(8/\delta)}{2n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(8/\delta)}{2n_{u}}}\ ,$
under the class shift assumption and conditions $\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}}(\bm{x})]\geq\gamma$ and $\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}}(\bm{x})\geq\gamma n_{u}$ for constant $\gamma\in(0,1)$.
Proof.
For simplicity, we denote by $\eta_{k}(\bm{x})=\Pr_{(\bm{x}^{\prime},y)\in\mathcal{D}}[y=k|\bm{x}^{\prime}=\bm{x}]$ for $k\in[\kappa+1]$ and $\bm{x}\in\mathcal{X}$. It is easy to observe that $\mu_{\mathcal{B}_{i}}(\bm{x})\in(0,1)$ and $\eta_{\kappa+1}(\bm{x})\in[0,1]$ for $\bm{x}\in\mathcal{X}$, and this follows that, from the Hoeffding’s inequality
$\left|\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})]-\frac{1}{n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})\right|\leq\sqrt{\frac{\ln(2/\delta_{1})}{2n_{u}}}\ ,$ (27)
with probability at least $1-\delta_{1}$ for some $\delta_{1}\in(0,\frac{1}{4})$. We similarly have
$\left|\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}}(\bm{x})]-\frac{1}{n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}}(\bm{x})\right|\leq\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}}\ ,$ (28)
with probability at least $1-\delta_{2}$ for some $\delta_{2}\in(0,\frac{1}{4})$, and it holds that
$\left|\mathbb{E}_{\mathcal{D}_{\text{kc}}}[\mu_{\mathcal{B}_{i}}(\bm{x})]-\frac{1}{n_{l}}\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\right|\leq\sqrt{\frac{\ln(2/\delta_{3})}{2n_{l}}}\ ,$ (29)
with probability at least $1-\delta_{3}$ for some $\delta_{3}\in(0,\frac{1}{4})$. From $\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\in[0,1]$, we have
$\left|\mathbb{E}_{\mathcal{D}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right]-\frac{1}{n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right|\leq\sqrt{\frac{\ln(2/\delta_{4})}{2n_{u}}}\ ,$ (30)
with probability at least $1-\delta_{4}$ for some $\delta_{4}\in(0,\frac{1}{4})$. From the class shift assumption in Eqn. (1), we have
$\mathbb{E}_{\mathcal{D}_{\text{kc}}}[\mu_{\mathcal{B}_{i}}(\bm{x})]=\mathbb{E}_{\mathcal{D}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})|y\in\left[\kappa\right]\right]$
$=\leavevmode\nobreak\ \frac{1}{\Pr_{\mathcal{D}}\left[y\in[\kappa]\right]}\mathbb{E}_{\mathcal{D}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right]=\frac{1}{1-\theta}\mathbb{E}_{\mathcal{D}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right]\ ,$
and we further have, from Eqn. (30)
$\left|\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]-\frac{1}{(1-\theta)n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right|\leq\frac{1}{1-\theta}\sqrt{\frac{\ln(2/\delta_{4})}{2n_{u}}}\ ,$ (31)
with probability at least $1-\delta_{4}$. For simplicity, we denote by
$\hat{\theta}_{\mathcal{B}_{i}}=\frac{\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})}{\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})}\ .$ (32)
From $\mu_{\mathcal{B}_{i}}(\bm{x})=\Pr[\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]$, we have
$\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]=\frac{\Pr_{\mathcal{D}}[y=\kappa+1,\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]}{\Pr_{\mathcal{D}}[\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]}=\frac{\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})]}{\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]}\ ,$
and this follows that, from Eqn. (32)
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\hat{\theta}_{\mathcal{B}_{i}}\right|=\left|\frac{\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})]}{\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]}-\frac{\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})}{\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})}\right|\ ,$
and we further have, from $\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\geq\gamma$ and $\frac{1}{n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\geq\gamma$
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\hat{\theta}_{\mathcal{B}_{i}}\right|$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{\frac{1}{n_{u}}\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})}\left|\frac{1}{n_{u}}\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})]\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})-\frac{1}{n_{u}}\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{\gamma^{2}}\left|\frac{1}{n_{u}}\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})]\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})-\frac{1}{n_{u}}\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{\gamma^{2}}\left|\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})]-\frac{1}{n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})\right|+\frac{1}{\gamma^{2}}\left|\frac{1}{n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})-\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\right|\ ,$ (33)
where the last inequality holds from Lemma 3, $\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})\eta_{\kappa+1}(\bm{x})]\in[0,1]$ and $\mathbb{E}_{\mathcal{D}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\in[0,1]$. From Eqns. (27)-(28) and (33), we could obtain that
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\hat{\theta}_{\mathcal{B}_{i}}\right|\leq\frac{1}{\gamma^{2}}\left(\sqrt{\frac{\ln(2/\delta_{1})}{2n_{u}}}+\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}}\right)\ ,$ (34)
with probability at least $1-\delta_{1}-\delta_{2}$. On the other hand, we have, from Eqn. (32) and $\sum_{k\in[\kappa+1]}\eta_{k}(\bm{x})=1$ for $\bm{x}\in\mathcal{X}$
$\left|\hat{\theta}_{\mathcal{B}_{i}}-\left(1-\sum_{(\bm{x},y)\in S_{l}}\frac{(1-\theta)n_{u}\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{l}n_{\mathcal{B}_{i},u}}\right)\right|=\left|\sum_{(\bm{x},y)\in S_{l}}\frac{(1-\theta)n_{u}\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{l}n_{\mathcal{B}_{i},u}}-\frac{\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})}{\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})}\right|\ ,$
and this follows that, from $\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}}(\bm{x})\geq\gamma n_{u}$ and $n_{\mathcal{B}_{i},u}={\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}}(\bm{x})}$
$\left|\hat{\theta}_{\mathcal{B}_{i}}-\left(1-\sum_{(\bm{x},y)\in S_{l}}\frac{(1-\theta)n_{u}\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{l}n_{\mathcal{B}_{i},u}}\right)\right|$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{(1-\theta)n_{u}}{{\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})}}\left|\frac{1}{n_{l}}\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})-\frac{1}{(1-\theta)n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1-\theta}{\gamma}\left|\frac{1}{n_{l}}\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})-\frac{1}{(1-\theta)n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1-\theta}{\gamma}\left(\left|\frac{1}{n_{l}}\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})-\mathbb{E}_{\mathcal{D}_{\text{kc}}}[\mu_{\mathcal{B}_{i}}(\bm{x})]\right|+\left|\mathbb{E}_{\mathcal{D}_{\text{kc}}}[\mu_{\mathcal{B}_{i}}(\bm{x})]-\frac{1}{(1-\theta)n_{u}}\sum_{\bm{x}\in S_{u}}\mu_{\mathcal{B}_{i}}(\bm{x})\sum_{k\in[\kappa]}\eta_{k}(\bm{x})\right|\right)\ .$ (35)
From Eqns. (29), (31) and (35), we have
$\left|\hat{\theta}_{\mathcal{B}_{i}}-\left(1-\sum_{(\bm{x},y)\in S_{l}}\frac{(1-\theta)n_{u}\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{l}n_{\mathcal{B}_{i},u}}\right)\right|\leq\frac{1-\theta}{\gamma}\sqrt{\frac{\ln(2/\delta_{3})}{2n_{l}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(2/\delta_{4})}{2n_{u}}}\ ,$ (36)
with probability at least $1-\delta_{3}-\delta_{4}$. We finally have, from Definition 2
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\left(1-\sum_{(\bm{x},y)\in S_{l}}\frac{(1-\theta)n_{u}\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{l}n_{\mathcal{B}_{i},u}}\right)\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\hat{\theta}_{\mathcal{B}_{i}}\right|+\left|\hat{\theta}_{\mathcal{B}_{i}}-\left(1-\sum_{(\bm{x},y)\in S_{l}}\frac{(1-\theta)n_{u}\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{l}n_{\mathcal{B}_{i},u}}\right)\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{\gamma^{2}}\sqrt{\frac{\ln(2/\delta_{1})}{2n_{u}}}+\frac{1}{\gamma^{2}}\sqrt{\frac{\ln(2/\delta_{2})}{2n_{u}}}+\frac{1-\theta}{\gamma}\sqrt{\frac{\ln(2/\delta_{3})}{2n_{l}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(2/\delta_{4})}{2n_{u}}}\ ,$
with probability at least $1-\sum_{i\in[4]}\delta_{i}$, where the last inequality holds from Eqns. (34) and (36). This completes the proof by setting $\delta_{1}=\delta_{2}=\delta_{3}=\delta_{4}=\delta/4$. ∎
Proof of Lemma 2. From Lemma 6, Lemma 2 holds for $k=\kappa+1$, and it remains to consider the case of $k\in[\kappa]$. It is easy to observe that $\mu_{\mathcal{B}}(\bm{x})\in(0,1)$ for any $\bm{x}\in\mathcal{X}$, and it follows that, from the Hoeffding’s inequality
$\left|\frac{1}{n_{l}}\sum\limits_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})-\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]\right|\leq\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}\ ,$ (37)
with probability at least $1-\delta_{1}$ for some $\delta_{1}\in(0,\frac{1}{3})$. We similarly have, with probability at least $1-\delta_{1}$
$\left|\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right]-\frac{1}{n_{l}}\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right|\leq\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}\ .$ (38)
From that $\mathcal{D}_{\text{kc}}$ is the marginal distribution of $\mathcal{D}$ over $\mathcal{X}\times[\kappa]$ and $\mu_{\mathcal{B}_{i}}(\bm{x})=\Pr[\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]$, we have
$\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|y\in[\kappa],\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]=\Pr_{(\bm{x},y)\sim\mathcal{D}_{\text{kc}}}[y=k|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]=\frac{\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right]}{\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]}\ ,$
and it follows that, from some simple calculations
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|y\in[\kappa],\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\frac{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]}{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})}\right|$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \left|\frac{\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right]}{\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]}-\frac{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]}{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})}\right|$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\left|\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right]\sum\limits_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})-\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right|}{\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]\sum\limits_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})}\ .$ (39)
From Eqn. (39), $\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}}(\bm{x})\right]\geq\gamma$ and $\sum\limits_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}}(\bm{x})\geq\gamma n_{l}$, we have
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|y\in[\kappa],\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\frac{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]}{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})}\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{\gamma^{2}}\left|\frac{1}{n_{l}}\sum\limits_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right]-\frac{1}{n_{l}}\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{\gamma^{2}}\left|\frac{1}{n_{l}}\sum\limits_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})-\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\right]\right|+\frac{1}{\gamma^{2}}\left|\mathbb{E}_{\mathcal{D}_{\text{kc}}}\left[\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right]-\frac{1}{n_{l}}\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right|\ ,$
where the last inequality holds from Lemma 3, and we further have, from Eqns. (37)-(38)
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|y\in[\kappa],\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\frac{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]}{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})}\right|\leq\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}\ ,$ (40)
with probability at least $1-2\delta_{1}$. From Lemma 6, we have
$\left|\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]\right|\leq\frac{1-\theta}{\gamma}\sqrt{\frac{\ln(8/\delta_{2})}{2n_{l}}}+\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(8/\delta_{2})}{2n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(8/\delta_{2})}{2n_{u}}}\ ,$ (41)
with probability at least $1-\delta_{2}$ for some $\delta_{2}\in(0,\frac{1}{3})$. On the other hand, we also have
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|y\in[\kappa],\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]\Pr_{(\bm{x},y)\sim\mathcal{D}}[y\in[\kappa]|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\frac{1-\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})}{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})}\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]\right|$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|y\in[\kappa],\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\frac{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})\mathbb{I}[y=k]}{\sum_{(\bm{x},y)\in S_{l}}\mu_{\mathcal{B}_{i}}(\bm{x})}\right|$
$\leavevmode\nobreak\ +\left|\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=\kappa+1|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]\right|\ ,$
and it follows that, with probability at least $1-2\delta_{1}-\delta_{2}$, from Eqns. (40)-(41)
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})\right|\leq\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(2/\delta_{1})}{2n_{l}}}+\frac{1-\theta}{\gamma}\sqrt{\frac{\ln(8/\delta_{2})}{2n_{l}}}+\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(8/\delta_{2})}{2n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(8/\delta_{2})}{2n_{u}}}\ ,$
and we further have, by setting $\delta_{1}=\delta_{2}=\delta/3$
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})\right|\leq\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(6/\delta)}{2n_{l}}}+\frac{1-\theta}{\gamma}\sqrt{\frac{\ln(24/\delta)}{2n_{l}}}+\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(24/\delta)}{2n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(24/\delta)}{2n_{u}}}\ ,$
with probability at least $1-\delta$, which completes the proof. ∎
A.4 Proof of Theorem 2
We begin with a useful lemma as follows.
Lemma 7.
For $i\in[2t+2]\setminus[t]$ and $\mathcal{L}^{*}_{\mathcal{B}_{i}}$ defined by Eqn. (9), we have
$\mathcal{L}^{*}_{\mathcal{B}_{i}}=1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]\right)^{2}\ .$
Proof.
For any $\bm{w}\in\Delta_{\kappa+1}$, we have, by some simple calculations
$E_{\mathcal{D}}\left[\|\bm{w}-\tilde{\bm{y}}\|^{2}_{2}|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})\right]$
$=\leavevmode\nobreak\ \sum_{k\in[\kappa+1]}\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]\left(\left(1-w_{k}\right)^{2}-w^{2}_{k}+\sum_{t\in[\kappa+1]}w^{2}_{t}\right)$
= $=\leavevmode\nobreak\ 1+\sum_{k\in[\kappa+1]}\left(w_{k}-\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]\right)^{2}-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]\right)^{2}$
$\geq\leavevmode\nobreak\ 1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]\right)^{2}\ ,$
and this follows that, from Eqn. (9)
$\mathcal{L}^{*}_{\mathcal{B}_{i}}=\min_{\bm{w}\in\Delta_{\kappa+1}}E_{\mathcal{D}}\left[\|\bm{w}-\tilde{\bm{y}}\|^{2}_{2}|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})\right]\geq 1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]\right)^{2}\ .$ (42)
Let $w^{\prime}_{k}=\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]$ for $k\in[\kappa+1]$, and we have
$E_{\mathcal{D}}\left[\mathcal{L}(\bm{w}^{\prime},\tilde{\bm{y}})|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})\right]=1-\sum_{k\in[\kappa+1]}\left(\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\to\mathcal{B}_{i})]\right)^{2}\geq\mathcal{L}^{*}_{\mathcal{B}_{i}}\ ,$
which completes the proof by combining it with Eqn. (42). ∎
Proof of Theorem 2. From Lemma 2 and Lemma 6, and we have
$\left|\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]-\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})\right|\leq\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(6/\delta)}{2n_{l}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(24/\delta)}{2n_{l}}}+\frac{2}{\gamma^{2}}\sqrt{\frac{\ln(24/\delta)}{2n_{u}}}+\frac{1}{\gamma}\sqrt{\frac{\ln(24/\delta)}{2n_{u}}}\ ,$ (43)
for $k\in[\kappa+1]$ with probability at least $1-\delta_{1}$ for some $\delta_{1}\in(0,\frac{1}{\kappa+1})$. For simplicity, we denote by
$\theta_{\mathcal{B}_{i},k}=\Pr_{(\bm{x},y)\sim\mathcal{D}}[y=k|\mathcal{I}(\bm{x}\rightarrow\mathcal{B}_{i})]\ ,$
for $k\in[\kappa+1]$. From Eqn. (8), we have
$\left|\mathcal{L}^{*}_{\mathcal{B}_{i}}-\mathcal{G}_{\mathcal{B}_{i}}(S_{l},S_{u})\right|$$\leq$ $\sum_{k\in[\kappa+1]}\left|\vartheta^{2}_{\mathcal{B}_{i},k}(S_{l},S_{u})-\theta^{2}_{\mathcal{B}_{i},k}\right|$
$\leq$ $\sum_{k\in[\kappa+1]}\left|\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})+\theta_{\mathcal{B}_{i},k}\right|\times\left|\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})-\theta_{\mathcal{B}_{i},k}\right|$
$\leq$ $\sum_{k\in[\kappa+1]}2\left|\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})-\theta_{\mathcal{B}_{i},k}\right|\ ,$
where the last inequality holds from $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})\in[0,1]$ and $\theta_{\mathcal{B}_{i},k}\in[0,1]$, and this follows that from Eqn. (43)
$\left|\mathcal{L}^{*}_{\mathcal{B}_{i}}-\mathcal{G}_{\mathcal{B}_{i}}(S_{l},S_{u})\right|$
$\leq\leavevmode\nobreak\ \frac{4(\kappa+1)}{\gamma^{2}}\sqrt{\frac{\ln(6/\delta_{1})}{2n_{l}}}+\frac{2(\kappa+1)}{\gamma}\sqrt{\frac{\ln(24/\delta_{1})}{2n_{l}}}+\frac{4(\kappa+1)}{\gamma^{2}}\sqrt{\frac{\ln(24/\delta_{1})}{2n_{u}}}+\frac{2(\kappa+1)}{\gamma}\sqrt{\frac{\ln(24/\delta_{1})}{2n_{u}}}\ ,$
with probability at least $1-(\kappa+1)\delta_{1}$, and this completes the proof by setting $\delta_{1}=\delta/(\kappa+1)$. ∎
Datasets Our LACForest GLAC EULAC EVM PAC-iForest OSNN LACU-SVM OVR-SVM
segment .9380$\pm$.0196 .8626$\pm$.0570$\bullet$ .9102$\pm$.0384$\bullet$ .8450$\pm$.1321$\bullet$ .6598$\pm$.0825$\bullet$ .6478$\pm$.0361$\bullet$ .6654$\pm$.0763$\bullet$ .6920$\pm$.0677$\bullet$
texture .9019$\pm$.0190 .9026$\pm$.0448 .8989$\pm$.0310 .8682$\pm$.0267$\bullet$ .6975$\pm$.0643$\bullet$ .6948$\pm$.0229$\bullet$ .7218$\pm$.0299$\bullet$ .7072$\pm$.0383$\bullet$
optdigits .9269$\pm$.0266 .8863$\pm$.0302$\bullet$ .9291$\pm$.0248 .9041$\pm$.0227$\bullet$ .7231$\pm$.0461$\bullet$ .7245$\pm$.0170$\bullet$ .7561$\pm$.0318$\bullet$ .8002$\pm$.0382$\bullet$
satimage .8478$\pm$.0415 .7361$\pm$.1022$\bullet$ .8249$\pm$.0552$\bullet$ .7090$\pm$.0714$\bullet$ .6888$\pm$.0610$\bullet$ .5402$\pm$.0437$\bullet$ .5960$\pm$.0247$\bullet$ .5423$\pm$.0641$\bullet$
landset .9083$\pm$.0330 .8435$\pm$.0322$\bullet$ .8619$\pm$.0422$\bullet$ .7921$\pm$.0520$\bullet$ .7223$\pm$.0767$\bullet$ .5685$\pm$.0395$\bullet$ .6219$\pm$.0178$\bullet$ .5741$\pm$.0493$\bullet$
mfcc .8806$\pm$.1015 .7205$\pm$.1277$\bullet$ .8832$\pm$.0892 .8541$\pm$.1225$\bullet$ .6186$\pm$.1372$\bullet$ .4533$\pm$.0603$\bullet$ .4849$\pm$.0507$\bullet$ .6346$\pm$.0260$\bullet$
usps .8696$\pm$.0276 .8549$\pm$.0339$\bullet$ .8702$\pm$.0322 .7800$\pm$.0654$\bullet$ .5538$\pm$.0792$\bullet$ .6940$\pm$.0225$\bullet$ .7089$\pm$.0482$\bullet$ .7788$\pm$.0343$\bullet$
har .8972$\pm$.0310 .8885$\pm$.0351$\bullet$ .8844$\pm$.0289$\bullet$ .4526$\pm$.0701$\bullet$ .5502$\pm$.0688$\bullet$ .5510$\pm$.0443$\bullet$ .4879$\pm$.0325$\bullet$ .5426$\pm$.0343$\bullet$
mapping .6477$\pm$.1041 .6021$\pm$.0953$\bullet$ .6777$\pm$.0975$\circ$ .5962$\pm$.1690$\bullet$ .5083$\pm$.1406$\bullet$ .5198$\pm$.0870$\bullet$ .3583$\pm$.1329$\bullet$ .4090$\pm$.1371$\bullet$
pendigits .9224$\pm$.0238 .8729$\pm$.0297$\bullet$ .9146$\pm$.0322$\bullet$ .8834$\pm$.0347$\bullet$ .7523$\pm$.0849$\bullet$ .6581$\pm$.0274$\bullet$ .7865$\pm$.0467$\bullet$ .7114$\pm$.0339$\bullet$
drybean .8814$\pm$.0252 .8821$\pm$.0742 .8906$\pm$.0264$\circ$ .7835$\pm$.0557$\bullet$ .6768$\pm$.0885$\bullet$ .6477$\pm$.0545$\bullet$ .7397$\pm$.0416$\bullet$ .6346$\pm$.0619$\bullet$
letter .6870$\pm$.0437 .4820$\pm$.0902$\bullet$ .5479$\pm$.0807$\bullet$ .6728$\pm$.0340$\bullet$ .4366$\pm$.0533$\bullet$ .6091$\pm$.0281$\bullet$ .4918$\pm$.0558$\bullet$ .5134$\pm$.0658$\bullet$
shuttle .7765$\pm$.1906 .7057$\pm$.1027$\bullet$ .8048$\pm$.1567$\circ$ .4576$\pm$.0729$\bullet$ .3577$\pm$.0468$\bullet$ .3618$\pm$.0880$\bullet$ .4217$\pm$.1264$\bullet$ .3843$\pm$.1032$\bullet$
drive .8466$\pm$.0519 .5711$\pm$.1358$\bullet$ .7526$\pm$.0806$\bullet$ .7660$\pm$.0434$\bullet$ .3383$\pm$.0735$\bullet$ .6433$\pm$.0208$\bullet$ .3266$\pm$.0615$\bullet$ .2832$\pm$.1169$\bullet$
senseveh .7962$\pm$.0254 .7418$\pm$.0375$\bullet$ .7752$\pm$.0194$\bullet$ .5464$\pm$.0552$\bullet$ .5036$\pm$.1359$\bullet$ .5349$\pm$.0437$\bullet$ .5726$\pm$.0398$\bullet$ .5548$\pm$.0574$\bullet$
mnist .8224$\pm$.0392 .7885$\pm$.0317$\bullet$ .8114$\pm$.0450$\bullet$ .4030$\pm$.0888$\bullet$ .5282$\pm$.0840$\bullet$ .6932$\pm$.0283$\bullet$ .6749$\pm$.0504$\bullet$ .7555$\pm$.0310$\bullet$
fmnist .7645$\pm$.0324 .7465$\pm$.0343$\bullet$ .7120$\pm$.0587$\bullet$ .5524$\pm$.0723$\bullet$ .5704$\pm$.0620$\bullet$ .5679$\pm$.0612$\bullet$ .6287$\pm$.0609$\bullet$ .6039$\pm$.0709$\bullet$
average .8421$\pm$.0813 .7699$\pm$.1212 .8206$\pm$.0986 .6980$\pm$.1608 .5815$\pm$.1229 .5947$\pm$.0934 .5908$\pm$.1368 .5954$\pm$.1395
win/tie/loss 16/2/0 10/3/4 17/0/0 17/0/0 17/0/0 17/0/0 17/0/0
Table 4: Experimental comparisons of macro-F1 score (mean±$\pm$±std) for compared methods, and $\bullet$/$\circ$ indicates that our approach is significantly better/worse than the corresponding method (paired $t$-test at $95\%$ significance level).
Appendix B Experimental Details
B.1 Evaluation of our LACForest approach
Datasets. All of datasets could be downloaded from https://archive.ics.uci.edu/datasets and https://www.openml.org, and we normalize the features of raw data into the range of $[0,1]$.
Parameter settings. For our LACForest, We set the random trees number $m=100$, $\gamma=0.01$ and $\tau=\lfloor\sqrt{d}\rfloor$ on all datasets, where $d$ denotes the data dimension. Parameters for other methods are set according to their respective references.
Performance measure. Following previous works Zhang et al. [2020]; Shu et al. [2023], we use three measures to evaluate the performance of compared methods, including accuracy, macro-F1 and AUC on the detection of augmented classes. For simplicity, we denote by $S_{t}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{t}}$ the testing dataset and $f:\mathcal{X}\rightarrow\mathcal{Y}$ the learned model. We present the performance measures in details as follows:
  • Accuracy: the average predictive accuracy on testing data, i.e.,
    $\text{ACC}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\mathbb{I}\left[f(\bm{x}_{i})=y_{i}\right]\ .$
  • Macro-F1: the average F1 score on testing data, i.e.,
[…]
    where $P_{k}$ and $R_{k}$ stand for the precision and recall of the $k$-th class, respectively.
  • AUC: the average AUC score on identifying augmented data, i.e.,
    $\text{AUC}=\frac{1}{|S^{+}_{t}|\times|S^{-}_{t}|}\sum_{\bm{x}^{+}\in S^{+}_{t}}\sum_{\bm{x}^{-}\in S^{-}_{t}}\mathbb{I}\left[f_{\kappa+1}(\bm{x}^{+})>f_{\kappa+1}(\bm{x}^{-})\right]\ ,$
    where $f_{\kappa+1}(\bm{x})$ stands for the augmented class score w.r.t. model $f$, and $S^{+}_{t}$ and $S^{-}_{t}$ could be defined as follows:
    $S^{+}_{t}=\{\bm{x}\colon(\bm{x},y)\in S_{t}\wedge y=\kappa+1\}\quad\text{ and }\quad S^{-}_{t}=\{\bm{x}\colon(\bm{x},y)\in S_{t}\wedge y\in[\kappa]\}\ .$
Additional experimental results. We present the average Macro-F1 scores and AUC scores over 100 different configurations for each dataset in Table 4 and Table 5, respectively. As can be seen, our approach exhibits better performance than other compared methods, since the win/tie/loss count clearly indicates that our approach wins in most times.
Datasets Our LACForest GLAC EULAC EVM PAC-iForest OSNN LACU-SVM OVR-SVM
segment .9891$\pm$.0082 .9687$\pm$.0141$\bullet$ .9807$\pm$.0141$\bullet$ .9423$\pm$.0309$\bullet$ .7547$\pm$.1005$\bullet$ .9326$\pm$.0447$\bullet$ .7019$\pm$.0254$\bullet$ .7631$\pm$.1040$\bullet$
texture .9799$\pm$.0083 .9842$\pm$.0088$\circ$ .9875$\pm$.0084$\circ$ .9546$\pm$.0179$\bullet$ .7925$\pm$.0460$\bullet$ .9054$\pm$.0292$\bullet$ .8207$\pm$.0588$\bullet$ .8133$\pm$.0904$\bullet$
optdigits .9894$\pm$.0072 .9798$\pm$.0075$\bullet$ .9916$\pm$.0054 .9741$\pm$.0101$\bullet$ .7948$\pm$.0484$\bullet$ .9609$\pm$.0142$\bullet$ .8567$\pm$.0312$\bullet$ .9530$\pm$.0227$\bullet$
satimage .9642$\pm$.0209 .9281$\pm$.0504$\bullet$ .9542$\pm$.0264$\bullet$ .8051$\pm$.0652$\bullet$ .8551$\pm$.0654$\bullet$ .7230$\pm$.1056$\bullet$ .7524$\pm$.0564$\bullet$ .7891$\pm$.0700$\bullet$
landset .9869$\pm$.0054 .9625$\pm$.0107$\bullet$ .9685$\pm$.0120$\bullet$ .9039$\pm$.0380$\bullet$ .8599$\pm$.0675$\bullet$ .8803$\pm$.0445$\bullet$ .8231$\pm$.0355$\bullet$ .8454$\pm$.0425$\bullet$
mfcc .9877$\pm$.0044 .9569$\pm$.0141$\bullet$ .9871$\pm$.0066 .9215$\pm$.0589$\bullet$ .8691$\pm$.0476$\bullet$ .9197$\pm$.0598$\bullet$ .8572$\pm$.0393$\bullet$ .8862$\pm$.0394$\bullet$
usps .9687$\pm$.0114 .9696$\pm$.0091 .9643$\pm$.0267$\bullet$ .8640$\pm$.0700$\bullet$ .6401$\pm$.1106$\bullet$ .8980$\pm$.0475$\bullet$ .8547$\pm$.0498$\bullet$ .9346$\pm$.0161$\bullet$
har .9662$\pm$.0210 .9740$\pm$.0146$\circ$ .9558$\pm$.0290$\bullet$ .5564$\pm$.0493$\bullet$ .6375$\pm$.0956$\bullet$ .5169$\pm$.0824$\bullet$ .5752$\pm$.0628$\bullet$ .6400$\pm$.1333$\bullet$
mapping .9482$\pm$.0157 .9093$\pm$.0322$\bullet$ .9486$\pm$.0154 .7786$\pm$.1671$\bullet$ .7564$\pm$.1336$\bullet$ .8202$\pm$.0897$\bullet$ .7659$\pm$.0894$\bullet$ .8192$\pm$.0771$\bullet$
pendigits .9796$\pm$.0105 .9692$\pm$.0115$\bullet$ .9820$\pm$.0211 .9475$\pm$.0236$\bullet$ .8498$\pm$.0694$\bullet$ .9437$\pm$.0285$\bullet$ .8872$\pm$.0436$\bullet$ .9209$\pm$.0254$\bullet$
drybean .9738$\pm$.0103 .9583$\pm$.0266$\bullet$ .9747$\pm$.0121 .8757$\pm$.0404$\bullet$ .8484$\pm$.0749$\bullet$ .8117$\pm$.0621$\bullet$ .6981$\pm$.0418$\bullet$ .7360$\pm$.1092$\bullet$
letter .8519$\pm$.0268 .7626$\pm$.0410$\bullet$ .8225$\pm$.0468$\bullet$ .8539$\pm$.0174 .6238$\pm$.0571$\bullet$ .7977$\pm$.0253$\bullet$ .7269$\pm$.0408$\bullet$ .6561$\pm$.0597$\bullet$
shuttle .9976$\pm$.0024 .9734$\pm$.0284$\bullet$ .9915$\pm$.0067$\bullet$ .9642$\pm$.0039$\bullet$ .9318$\pm$.0074$\bullet$ .9701$\pm$.0193$\bullet$ .8952$\pm$.0047$\bullet$ .8503$\pm$.0029$\bullet$
drive .9340$\pm$.0444 .7589$\pm$.0975$\bullet$ .8973$\pm$.0442$\bullet$ .8955$\pm$.0396$\bullet$ .6076$\pm$.0861$\bullet$ .8448$\pm$.0368$\bullet$ .6588$\pm$.0626$\bullet$ .5992$\pm$.1241$\bullet$
senseveh .8814$\pm$.0293 .8986$\pm$.0256$\circ$ .8958$\pm$.0268$\circ$ .6174$\pm$.0712$\bullet$ .6138$\pm$.1410$\bullet$ .6038$\pm$.0680$\bullet$ .6376$\pm$.0837$\bullet$ .6249$\pm$.0943$\bullet$
mnist .9435$\pm$.0292 .9341$\pm$.0248$\bullet$ .9330$\pm$.0330$\bullet$ .6189$\pm$.0477$\bullet$ .6414$\pm$.1197$\bullet$ .8370$\pm$.0297$\bullet$ .7989$\pm$.0456$\bullet$ .8632$\pm$.0337$\bullet$
fmnist .9216$\pm$.0299 .9020$\pm$.0379$\bullet$ .9097$\pm$.0330$\bullet$ .6725$\pm$.0409$\bullet$ .7132$\pm$.0942$\bullet$ .6755$\pm$.0658$\bullet$ .6425$\pm$.0531$\bullet$ .7016$\pm$.1201$\bullet$
average .9567$\pm$.0391 .9288$\pm$.0668 .9497$\pm$.0443 .8321$\pm$.1318 .7521$\pm$.1045 .8260$\pm$.1253 .7619$\pm$.0948 .7880$\pm$.1092
win/tie/loss 13/1/3 10/5/2 16/1/0 17/0/0 17/0/0 17/0/0 17/0/0
Table 5: Experimental comparisons of AUC (mean±$\pm$±std) on the detection of augmented class, and $\bullet$/$\circ$ indicates that our approach is significantly better/worse than the corresponding method (paired $t$-test at $95\%$ significance level).
B.2 Evaluation of our deep neural LACForest
Datasets. We follow the data setting of each dataset in previous studies Zhang et al. [2020]; Shu et al. [2023]. All the datasets contain ten classes, among which we randomly select six classes as known classes and four other classes as the augmented class. The detailed description of each dataset is presented as follows.
  • Mnist contains 28$\times$28 monochrome images of digital numbers. The labeled data contains instances of six known classes with 4000 instances per class. On the other hand, the unlabeled data contains unlabeled instances from 10 classes with 1000 instances in each class. Thus, the total number of training instances is 34000. The testing set consists of ten classes with 100 instances in each class, and thus the total number is 1000. The dataset can be downloaded from http://yann.lecun.com/exdb/mnist/.
  • Fmnist includes 28$\times$28 monochrome images of fashion apparel, and it takes similar data setting to that of mnist. The dataset can be downloaded from https://github.com/zalandoresearch/fashion-mnist.
  • Kuzushiji contains 28$\times$28 monochrome images of handwritten japanese characters, and it takes similar data setting to that of mnist. The dataset can be downloaded from https://github.com/rois-codh/kmnist.
  • Svhn is short for Street View House Numbers, which includes 32$\times$32 colored image samples of numbers. The training set consists of six known classes with 4500 subsampled instances in each class and ten unlabeled classes with 3000 instances in each class, and the total number of instances is 57000. The testing set contains instances of all ten classes with 100 instances in each class. SVHN dataset can be downloaded from http://ufldl.stanford.edu/housenumbers/.
  • Cifar10 contains natural images of size 32$\times$32. The training set contains six known classes with 5000 subsampled instances in each class and unlabeled instances from all ten classes with 900 subsampled instances in each class; thus, the total number of training instance is 39000. The testing set contains all ten classes with 100 instances per class. Cifar-10 dataset can be downloaded from https://www.cs.toronto.edu/~kriz/cifar.html.
Compared methods. This section introduces the details of compared approaches, where Deep-GLAC Shu et al. [2023] and Deep-EULAC Zhang et al. [2020] could be seen as direct extensions of GLAC and EULAC in section 5.1, respectively.
  • Softmax-T Hendrycks and Gimpel [2016] trains $\kappa$ classifiers $f_{1}(\bm{x}),\cdots,f_{\kappa}(\bm{x})$ to approximate the posterior probabilities $\Pr[y=k|\bm{x}]$ for $k\in[\kappa+1]$. An instance $\bm{x}$ is predicted to be the augmented class when the maximum outputs $\max_{k\in[\kappa]}f_{k}(\bm{x})$ is less than a given threshold.
  • Openmax Bendale and Boult [2016] can be seen as a calibrated version of Softmax-T, where replaces the Softmax layer with a new Openmax layer based on Weibull calibration.
  • G-Openmax Ge et al. [2017] is an extension of Openmax approach with generative neural networks.
  • OSRCI Neal et al. [2018] is short for Open-set Recognition using Counterfactual Images, which utilizes a counterfactual image generation technique to train an additional classifier for the augmented class.
  • ARPL Chen et al. [2021] is short for Adversarial Reciprocal Points Learning, which introduces a novel framework for open set recognition by leveraging reciprocal points to model latent open space.
Parameter settings. For our approach, we construct 3 neural trees of depth $l=6$, and set the hyper-parameter $\lambda_{\text{ce}}=1$ for all datasets. We set the number of epochs as 500 for all datasets, and take a relatively large batch size 512 for both labeled and unlabeled data according to Theorem 2. The loss is optimized by SGD optimizer with weight decay $5\times 10^{-3}$. The initial learning rate is set to $10^{-2}$ and is progressively reduced to $10^{-3}$ through a cosine annealing scheduler. For the estimation of proportion $\theta$, we randomly sample 1000 labeled examples and 1000 unlabeled instances for ten times, and take the average estimated proportion by previous method Ramaswamy et al. [2016]. For other compared methods, we select their parameters according to their corresponding references.
Additional experimental results. We present the average Macro-F1 scores and AUC scores for each dataset in Table 6 and Table 7, respectively. It is observable that our approach exhibits better performance than other compared methods in most times, which indicates the effectiveness of our approach under different performance measures.
Datasets Our approach Deep-GLAC Deep-EULAC ARPL G-Openmax OSRCI Openmax Softmax-T
mnist .9833$\pm$.0027 .9778$\pm$.0035 .9601$\pm$.0027 .9390$\pm$.0171 .8995$\pm$.0053 .9171$\pm$.0044 .8972$\pm$.0041 .8912$\pm$.0030
fmnist .8986$\pm$.0141 .8945$\pm$.0171 .8436$\pm$.0076 .7814$\pm$.0092 .6925$\pm$.0135 .7099$\pm$.0084 .6832$\pm$.0178 .6297$\pm$.0102
kuzushiji .9629$\pm$.0064 .8867$\pm$.0038 .9504$\pm$.0041 .9117$\pm$.0120 .8582$\pm$.0053 .8611$\pm$.0038 .8503$\pm$.0043 .8333$\pm$.0060
svhn .9237$\pm$.0140 .8936$\pm$.0145 .8236$\pm$.0094 .8050$\pm$.0080 .7871$\pm$.0157 .8039$\pm$.0206 .8026$\pm$.0066 .7642$\pm$.0034
cifar10 .7936$\pm$.0299 .7911$\pm$.0429 .7119$\pm$.0231 .7271$\pm$.0031 .6777$\pm$.0205 .7061$\pm$.0133 .6957$\pm$.0391 .6857$\pm$.0264
average .9124$\pm$.0663 .8887$\pm$.0592 .8579$\pm$.0913 .8328$\pm$.0801 .7830$\pm$.0878 .7996$\pm$.0829 .7858$\pm$.0843 .7608$\pm$.0950
Table 6: Experimental comparisons of macro-F1 score (mean±$\pm$±std) over 5 image datasets, and the best performance is highlighted in bold.
Datasets Our approach Deep-GLAC Deep-EULAC ARPL G-Openmax OSRCI Openmax Softmax-T
mnist .9983$\pm$.0019 .9974$\pm$.0010 .9927$\pm$.0009 .9517$\pm$.0452 .9396$\pm$.0031 .9407$\pm$.0016 .9399$\pm$.0032 .9367$\pm$.0029
fmnist .9784$\pm$.0061 .9751$\pm$.0149 .9505$\pm$.0080 .8607$\pm$.0046 .8423$\pm$.0042 .8481$\pm$.0049 .8397$\pm$.0070 .8292$\pm$.0042
kuzushiji .9939$\pm$.0026 .9927$\pm$.0035 .9805$\pm$.0019 .9581$\pm$.0078 .9090$\pm$.0029 .9110$\pm$.0043 .9066$\pm$.0030 .8984$\pm$.0028
svhn .9795$\pm$.0020 .9839$\pm$.0051 .9444$\pm$.0073 .9195$\pm$.0026 .8678$\pm$.0059 .8969$\pm$.0029 .8654$\pm$.0055 .8535$\pm$.0063
cifar10 .9292$\pm$.0095 .9254$\pm$.0071 .8535$\pm$.0153 .8182$\pm$.0051 .7574$\pm$.0052 .7947$\pm$.0062 .7740$\pm$.0143 .7292$\pm$.0188
average .9759$\pm$.0246 .9749$\pm$.0259 .9443$\pm$.0489 .9016$\pm$.0541 .8632$\pm$.0626 .8783$\pm$.0514 .8651$\pm$.0570 .8494$\pm$.0706
Table 7: Experimental comparisons of AUC (mean±$\pm$±std) on the detection of augmented class over 5 image datasets. Notice that the best performance is highlighted in bold.